### v2

Abstract
We analyze Pareto optimality and competitive equilibria in a risk-exchange economy, where either all agents are risk seeking in an expected utility model, or they exhibit local risk-seeking behaviour in a rank-dependent utility model. A novel mathematical tool, the counter-monotonic improvement theorem, states that for any nonnegative allocation of the aggregate random payoff, there exists a counter-monotonic random vector, called a jackpot allocation, that is componentwise riskier than the original allocation, and thus preferred by risk-seeking agents. This result allows us to characterize Pareto optimality, the utility possibility frontier, and competitive equilibria with risk-seeking expected utility agents, and prove the first and second fundamental theorems of welfare economics in this setting. For rank-dependent utility agents that are neither risk averse or risk seeking, we show that jackpot allocations can be Pareto optimal for small-scale payoffs, but for large-scale payoffs they are dominated by proportional allocations, thus explaining the often-observed small-stake gambling behaviour in a risk sharing context. Such jackpot allocations are also equilibrium allocations for small-scale payoffs when there is no aggregate uncertainty.
Keywords: Gambling behaviour, counter-monotonicity, competitive equilibrium, convex utility functions, rank-dependent utility
1 Introduction
The exchange of risk is a fundamental issue in microeconomics, with various popular forms such as financial markets, insurance, gambling, and business cooperation. We consider the classic framework of a one-period risk exchange economy where an aggregate wealth (payoff) is distributed among a group of agents. Uncertainty in the aggregate monetary payoff is modelled by a random variable, which realizes at the end of the period, time $1$. A sharing plan, called an allocation, is represented by a random vector, designed at time 00 and realizing at time $1$, with components summing up to the total payoff. We assume all agents assess their risks or utilities under an objectively agreed probability, as in the framework of decision under risk.
Two important concepts in this risk exchange economy are Pareto optimality and Arrow–Debreu competitive equilibria. Under the classic setting of risk-averse agents, these two concepts are connected by the two fundamental theorems of welfare economics, which together state that Pareto-optimal allocations and equilibrium allocations are equivalent under some standard conditions (Arrow and Debreu, 1954; Arrow, 1964; Radner, 1968), such as the assumption of risk-averse expected utility (EU) agents. While risk aversion is mathematically tractable and normatively appealing, risk-seeking behaviour is widely empirically observed. Especially, local risk-seeking behaviour for small-stake payoffs is prominent in human behaviour. Various decision models were developed to accommodate this phenomenon, from using a non-concave utility function by Friedman and Savage (1948) and Markowitz (1952) in the EU model, to the modern rank-dependent utility (RDU) model (Quiggin, 1982, 1993) and the prospect theory (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992). For instance, a key feature of the cumulative prospect theory of Tversky and Kahneman (1992) is that agents are risk seeking for small-probability gains.
Although the risk-seeking behaviour is empirically relevant and theoretically important, studies on risk sharing problems for risk-seeking, or partially risk-seeking, agents are scarce. We believe that a major reason is its mathematical challenges. For instance, in the EU model, risk aversion corresponds to concave utility functions, guaranteeing nice properties of the utility possibility set and the existence of competitive equilibria; these are not guaranteed for convex utility functions. The ambitious aim of this paper is to bridge this gap by studying risk sharing for two classes of agents. First, we study the standard risk-seeking EU agents to gain crucial insights. Then, we turn to the empirically more relevant case of RDU agents following the model of Tversky and Kahneman (1992) with a concave-convex probability weighting function and a concave utility function. Such agents exhibit risk-seeking behaviour for small-stake and small-probability payoffs, but they are risk averse for large-stake or large-probability payoffs.
It is helpful to first take a close look into the case of risk-averse agents. An important observation for risk-averse agents is that Pareto and equilibrium allocations are comonotonic, and in equilibrium, every agent’s random payoffs are increasing functions of the total wealth. Comonotonicity is an extreme form of positive dependence using the language of probability theory. This can be interpreted as agents being “on the same boat” when losses or gains occur, and a simple example is a proportional allocation. This is intuitively appealing and can be formalized by a mathematical result called the comonotonic improvement theorem of Landsberger and Meilijson (1994). The latter result states that for any given random payoff vector, there exists a comonotonic random vector whose components are less risky in the sense of Rothschild and Stiglitz (1970) (mathematically, smaller in convex order) than those of the original vector. This establishes the important intuition that risk-averse agents always prefer comonotonic allocations.
Given the strong relation between risk-averse agents and positively dependent Pareto-optimal allocations, one might wonder if a notion of negative dependence is connected to risk-seeking agents; that is, in equilibrium, the agents are “in opposite boats” when losses or gains occur. Such extreme negative dependence is represented by so-called counter-monotonicity. Unfortunately, although counter-monotonicity and comonotonicity are conceptually and mathematically symmetric in dimension two, they are quite different in dimension three or more, with counter-monotonicity being much more technically challenging to handle (see Puccetti and Wang, 2015 for a review). For a reader unfamiliar with dependence concepts, imagine two cars moving in the same direction (comonotonicity) or moving in opposite directions (counter-monotonicity). Next, add another car. While it is easy to imagine three cars moving in the same direction (pairwise), it is conceptually impossible to imagine them moving in opposite directions (pairwise).
Our main contributions address the above challenges, and we summarize them below. Based on recent developments in dependence modelling, we introduce jackpot and scapegoat allocations, which are counter-monotonic allocations that are normalized. These allocations resemble extreme forms of gambling as either “winner-takes-all” or “loser-loses-all” (drawing straws) lotteries. Parallel to the comonotonic improvement theorem, we establish the counter-monotonic improvement theorem (Theorem 1) in Section 3, which states that for any given random payoff vector with nonnegative components, there exists a jackpot allocation whose components are riskier than those of the original vector. An immediate consequence of this result is that all nonnegative Pareto-optimal allocations are jackpot allocations for a group of strictly risk-seeking agents.
The counter-monotonic improvement theorem leads to many advances in risk sharing. By restricting to nonnegative payoffs, we characterize all Pareto-optimal allocations (Theorem 2) for risk-seeking EU agents in Section 4. We also show that in the homogeneous case (i.e., all agents have the same preferences), the utility possibility frontier is a simplex (Theorem 3). Under the risk-seeking EU setting, the two fundamental welfare theorems are established (Theorem 4) in Section 5. In particular, equilibrium allocations are jackpot allocations. In the homogeneous case, all jackpot allocations are also equilibrium allocations (Theorem 5) and a competitive equilibrium always exists with a unique equilibrium price (Theorem 6). The case of RDU agents is much more technically challenging, which we discuss in Section 6, assuming homogeneous RDU agents with a concave-convex probability weighting function and a concave utility function that is locally linear. It turns out, perhaps surprisingly, that some natural jackpot allocations are Pareto-optimal in this case (Theorem 7), even though the agents are not risk-seeking. Moreover, under some specific assumptions, we show that for small-scale aggregate payoffs it is Pareto optimal to gamble, whereas for large-scale payoffs it is better, or even optimal, to share proportionally. We also obtain a competitive equilibrium under the assumption of no aggregate uncertainty on the small-scale total endowment (Proposition 5).
As far as we know, our paper is the first to provide explicit forms of Pareto optima and equilibria for risk-seeking EU agents and typical RDU agents, and to explain the empirically noted small-stake gambling behaviour in a risk sharing context. The problem addressed in the paper is a fundamental one, but the solutions are built on recent developments in mathematics and probability theory. Especially, a randomization technique, requiring an atomless probability space (formalized in Assumption ER), is crucial to all results. This demands a probabilistic formulation of the problem, different from the traditional framework of a finite-dimensional Euclidean space, usually assumed when studying the exchange of commodities. For risk-averse agents, the finite-dimensional space works perfectly fine, because no extra randomness is needed to reach optimality or equilibrium. For risk-seeking and more general agents, the extra randomness is essential for reaching Pareto optimality or equilibrium, because such agents may choose to introduce randomness in addition to their random payoffs, as they prefer gambles over payoffs with certainty.
The paper has five appendices. Appendix A presents a few technical lemmas that are useful in the proofs. Proofs of technical results are presented in Appendix B. Appendix C contains some technical background for counter-monotonicity. Appendix D presents some generalizations of our model assumptions on RDU agents. Appendix E discusses the existence of competitive equilibria for general EU agents for the interested reader.
1.1 Literature review
Risk sharing problems have a long history and many formulations; for a general reference to Pareto optimality and competitive equilibria, see Mas-Colell et al. (1995, Chapter 10). While the economics literature does not always employ the terms comonotonicity and comonotonic allocations, these concepts have long been a subject of interest in these fields. Of direct relevance is the case of economies with a constant aggregate endowment, where comonotonic allocations are constant across states. In this context, comonotonic allocations are sometimes called “no-betting” or “risk-free allocations.” See, for instance, Billot et al. (2000), Rigotti et al. (2008), and Strzalecki and Werner (2011) in the literature on risk sharing under heterogeneous beliefs and ambiguity, and the more recent Beissner et al. (2024) for two RDU agents. Chateauneuf et al. (2000) analyzed comonotonic allocations with aggregate risk when all agents are ambiguity-averse Choquet expected utility maximizers.
The study of risk sharing with risk-seeking agents is relatively scarce, but two papers are relevant to our work. With some concave preferences (allowing risk-seeking EU agents), Araujo et al. (2017) studied the existence of individually rational Pareto optimal allocations on the space of infinite sequences. They also showed that at optimum, at most one risk-seeking agent has strictly positive consumption in every state of the world. This corresponds to a counter-monotonic allocation among the risk-seeking agents.
Araujo et al. (2018) studied a finite-state exchange economy with both risk-averse and risk-seeking agents, under an assumption that the total endowment of risk-averse agents at one state is large enough compared to all other states, a competitive equilibrium exist. In contrast to the above two papers, we obtain a complete and explicit characterization of Pareto optimality and competitive equilibria for risk-seeking EU agents, as well as some Pareto optimal allocations for RDU preferences that are neither concave nor convex.
Risk-seeking behaviour has received increasing attention in the recent literature; for instance, Crinich et al. (2013) showed that risk-seeking agents are also naturally prudent, just like risk-averse agents, and Müller et al. (2017) designed a stochastic dominance relation allowing for comparing mildly risk-seeking agents. The experiments of Holt and Laury (2002) and Noussair et al. (2014) both estimated about 10% of the studied subjects to be risk seeking; these subjects not only exhibit a local risk-seeking behaviour, but also fit well with a convex utility function, hence a globally risk-seeking model. Moreover, the empirical study of Jullien and Salanié (2000) estimated that British racetrack bettors have a risk-seeking utility function if the EU model is assumed.
The technique of comonotonic improvement was first introduced by Landsberger and Meilijson (1994) and subsequently extended by Dana and Meilijson (2003), Ludkovski and Rüschendorf (2008) and Carlier et al. (2012). The concept of counter-monotonicity in dimension greater than two was formally studied by Dall’Aglio (1972) and further by Dhaene and Denuit (1999) and Cheung and Lo (2014). Puccetti and Wang (2015) gave a review on dependence concepts. Lauzier et al. (2023a) obtained a stochastic representation of counter-monotonicity, which is used in Section 3 to introduce jackpot and scapegoat allocations.
Outside the context of risk sharing, comonotonicity is a central concept in its own right in decision-making under uncertainty, as in the classic decision models of Yaari (1987); Schmeidler (1989). As the natural counterpart to comonotonicity, counter-monotonicity is also called anti-comonotonicity, but it received much less attention until recently; see Aouani et al. (2021) and Principi et al. (2023).
Counter-monotonicity also appears in the context of risk sharing for agents using quantiles (called the Value-at-Risk in finance), even though these agents are neither risk seeking nor risk averse. Embrechts et al. (2018) and Weber (2018) studied quantile-based risk sharing problems on probability spaces with Pareto-optimal counter-monotonic allocations, and Embrechts et al. (2020) studied the case of heterogeneous beliefs. Lauzier et al. (2023b) contained a risk sharing problem where the optimal allocations entail both positive and negative dependence.
The RDU model was introduced by Quiggin (1982), and it coincides with the cumulative prospect theory of Tversky and Kahneman (1992) when only modelling the gain of payoffs; that is, payoffs do not involve losses. The shape of the concave-convex probability weighting functions in Section 6 was proposed by Tversky and Kahneman (1992). For its parameter calibration, see Wu and Gonzalez (1996). We refer to Wakker (2010) for a general treatment of popular decision models and the relationship between them.
2 Model setting
We consider a one-period economy, where uncertainty is realized at the end of the period. Let $\mathcal{X}$ be a convex cone of random variables, which represent random monetary payoffs at the end of the period. All random variables are defined on a fixed probability space $(\Omega,\mathcal{F},\mathbb{P})$, with $\mathcal{F}$ representing the collection of all events and $\mathbb{E}$ representing the expectation under $\mathbb{P}$. For instance, the space $\mathcal{X}$ can be chosen as the space $L^{1}$ of integrable random variables. We treat almost surely equal random variables as identical, and we omit “almost surely” in most equalities unless we want it emphasized. [^2] Two random variables $X$ and $Y$ are almost surely equal if $\mathbb{P}(X=Y)=1$, and we will simply write $X=Y$. Let $n$ be a positive integer and write $[n]=\{1,\dots,n\}$. Let $\mathbb{R}_{+}=[0,\infty)$ and $\Delta_{n}$ be the standard simplex in $\mathbb{R}^{n}$, that is
$\Delta_{n}=\left\{(\theta_{1},\dots,\theta_{n})\in\mathbb{R}_{+}^{n}:\sum_{i=1}^{n}\theta_{i}=1\right\},$
and write $\Delta_{n}(v)=v\Delta_{n}=\{v\bm{\theta}:\bm{\theta}\in\Delta_{n}\}$ for $v>0$. Throughout, for a scalar $z$ and a vector $\mathbf{y}=(y_{1},\dots,y_{n})$, we write $z\mathbf{y}$ for the vector $(zy_{1},\dots,zy_{n})$. Denote by $\mathbf{0}$ and $\mathbf{1}$ the vectors $(0,\dots,0)$ and $(1,\dots,1)$ in $\mathbb{R}^{n}$, and thus $y\mathbf{1}=(y,\dots,y)$ and $\mathbf{1}/y=(1/y,\dots,1/y)$ for $y>0$. We will use boldface (capital) letters for (random) vectors.
Our setting of risk sharing concerns $n$ agents who share a random variable $X\in\mathcal{X}$, which is interpreted as a total random wealth to be allocated among the agents. We will refer to the elements of $\mathcal{X}$ as risks. We allow the risk $X$ to take both positive (meaning gains) and negative (meaning losses) values, although, in some later sections, we will restrict to positive or negative risks. We are mainly interested in the situation of $n\geq 3$. The set of all allocations of $X\in\mathcal{X}$ is defined as
$\mathbb{A}_{n}(X)=\left\{(X_{1},\dots,X_{n})\in\mathcal{X}^{n}:\sum_{i=1}^{n}X_{i}=X\right\}.$
That is, an allocation of $X$ is a random vector whose components sum to $X$. This means the wealth is completely redistributed among the agents without any transfers to or from outside the group. Note that the choice of $\mathcal{X}$ is important in the definition of $\mathbb{A}_{n}$ as it restricts the possible allocations.
For each $i\in[n]$, the preference $\succsim_{i}$ of agent $i$ is numerically represented by a preference functional $\mathcal{U}_{i}$, that is,
$X\succsim_{i}Y~{}\iff~{}\mathcal{U}_{i}(X)\geq\mathcal{U}_{i}(Y).$
The value $\mathcal{U}_{i}(X)$ is referred to as the utility of $X$ for agent $i$. We assume that each $\mathcal{U}_{i}$ represents a decision model under risk, and all agents agree on the probability measure $\mathbb{P}$. Formally, this means that if $X$ and $Y$ are equally distributed, denoted by $X\buildrel\mathrm{d}\over{=}Y$, then $\mathcal{U}_{i}(X)=\mathcal{U}_{i}(Y)$. For instance, $\mathcal{U}_{i}$ may be an EU preference functional $\mathcal{U}_{i}:X\mapsto\mathbb{E}[u_{i}(X)]$ for some increasing function $u_{i}:\mathbb{R}\to\mathbb{R}$ (called a utility function); such agents are called EU agents or EU maximizers. Later, we will also consider RDU agents (defined in Section 6). Throughout, terms like “increasing” and “decreasing” are in the non-strict sense.
We will study two types of optimality in risk sharing: Pareto optimality and Arrow-Debreu competitive equilibria, explained below. We also define a related concept of individual rationality.
Pareto optimality. For two allocations $\mathbf{X}=(X_{1},\dots,X_{n})$ and $\mathbf{Y}=(Y_{1},\dots,Y_{n})$ in $\mathbb{A}_{n}(X)$, we say that $\mathbf{X}$ dominates $\mathbf{Y}$ if $\mathcal{U}_{i}(X_{i})\geq\mathcal{U}_{i}(Y_{i})$ for all $i\in[n]$, and the domination is strict if at least one of the inequalities is strict. The allocation $\mathbf{X}$ is Pareto optimal if it is not strictly dominated by any allocation in $\mathbb{A}_{n}(X)$. Pareto optimality is a classic notion of optimality or efficiency, and it is closely connected to optimization of a linear combination of the utilities. For $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}_{+}^{n}\setminus\{\mathbf{0}\}$, [^3] The vector $\bm{\lambda}$ is usually called a Negishi weight vector (Negishi (1960)). we say that an allocation $\mathbf{X}$ is ${\bm{\lambda}}$-optimal in $\mathbb{A}_{n}(X)$ if $\sum_{i=1}^{n}\lambda_{i}\mathcal{U}_{i}(X_{i})$ is maximized over $\mathbb{A}_{n}(X)$. We use the term sum optimality for the case $\lambda_{i}=1$ for every $i\in[n]$. It is well-known and straightforward to check that ${\bm{\lambda}}$-optimality for ${\bm{\lambda}}$ with positive components implies Pareto optimality. The converse holds under some additional conditions. [^4] In the converse statement, ${\bm{\lambda}}$ may have some zero components, and some conditions on the utility possibility set are needed; see e.g., Chapter 16 of Mas-Colell et al. (1995).
Competitive equilibria. Suppose that each of the agents has an initial endowment, summarized by the vector $\bm{\xi}=(\xi_{1},\dots,\xi_{n})\in\mathbb{A}_{n}(X)$. Consider the individual optimization problem for agent $i\in[n]$:
to $\mbox{to maximize}~{}\mathcal{U}_{i}(X_{i})~{}~{}~{}~{}\mbox{over $X_{i}\in\mathcal{X}_{i}$}~{}~{}~{}~{}\mbox{subject to}~{}\mathbb{E}^{Q}[X_{i}]\leq\mathbb{E}^{Q}[\xi_{i}];$ (1)
where $Q$ is a probability measure representing a linear pricing mechanism, $\mathbb{E}^{Q}$ is the expected value under $Q$, and for each $i\in[n]$, $\mathcal{X}_{i}$ is a set of possible choices of the position $X_{i}$ for agent $i$. In Section 4, we will choose $\mathcal{X}_{i}$ to be the set of all random variables $Y$ satisfying $0\leq Y\leq X$, where $X$ is assumed to be nonnegative. The constraint $\mathbb{E}^{Q}[X_{i}]\leq\mathbb{E}^{Q}[\xi_{i}]$ is a budget condition, typically holding with equality. The tuple $(X_{1},\dots,X_{n},Q)$ is a competitive equilibrium if (a) individual optimality holds: $X_{i}$ solves (1) for each $i\in[n]$; and (b) market clearance holds: $\sum_{i=1}^{n}X_{i}=X$. In a competitive equilibrium, $(X_{1},\dots,X_{n})$ is called an equilibrium allocation, and $Q$ is called an equilibrium price. We also call $(X_{1},\dots,X_{n})$ an equilibrium allocation of $X$, without mentioning its initial endowments, if it solves (1) for some $(\xi_{1},\dots,\xi_{n})$.
Individual rationality. For an initial endowment vector $\bm{\xi}\in\mathbb{A}_{n}(X)$, an allocation $\mathbf{X}\in\mathbb{A}_{n}(X)$ is individually rational if it dominates $\bm{\xi}$; that is, participation in risk sharing is beneficial for each agent.
For the same initial endowment vector, an equilibrium allocation is always individually rational, because of individual optimality. The concepts of Pareto-optimal allocations and equilibrium allocations are intimately connected through the two fundamental theorems of welfare economics. Under certain conditions, the first welfare theorem states that any equilibrium allocation is Pareto optimal, and the second welfare theorem states that any Pareto-optimal allocation is an equilibrium allocation for some initial endowments and equilibrium price. In Section 5, we will establish welfare theorems in the context of risk-seeking agents.
3 Counter-monotonic improvement and jackpot allocations
In this section, we first review the concepts of convex order and risk aversion and then introduce the counter-monotonic improvement theorem and jackpot allocations, which serve as fundamental tools for analyzing risk sharing problems with risk-seeking agents. In this section, $\mathcal{X}=L^{1}$.
3.1 Convex order, risk aversion, and comonotonicity
We first describe an essential concept that characterizes risk aversion. A random variable $X$ is said to be smaller than a random variable $Y$ in convex order, denoted by $X\leq_{\rm{cx}}Y$, if $\mathbb{E}[\phi(X)]\leq\mathbb{E}[\phi(Y)]$ for every convex function $\phi:\mathbb{R}\to\mathbb{R}$ provided that both expectations exist (see Rüschendorf (2013) and Shaked and Shanthikumar (2007)). The order $X\leq_{\rm{cx}}Y$ means that $X$ is less risky than $Y$ in the sense of Rothschild and Stiglitz (1970). Notice that if $X\leq_{\rm{cx}}Y$, then $\mathbb{E}[X]=\mathbb{E}[Y]$, meaning that the relation of convex order compares random variables with the same mean; for this reason, $Y$ is also called a mean-preserving spread of $X$ in the literature. [^5] A related concept is second-order stochastic dominance, denoted by $X\leq_{\rm ssd}Y$, meaning that $\mathbb{E}[\psi(X)]\leq\mathbb{E}[\psi(Y)]$ for every increasing concave function $\psi:\mathbb{R}\to\mathbb{R}$ provided the expectations exist. The two concepts are connected by $X\leq_{\rm cx}Y\iff Y\leq_{\rm ssd}X\mbox{~{}and~{}}\mathbb{E}[X]=\mathbb{E}[Y].$ If a preference functional $\mathcal{U}$ satisfies
$X\leq_{\rm cx}Y~{}\Longrightarrow~{}\mathcal{U}(X)\geq\mathcal{U}(Y),$
then we say that $\mathcal{U}$ is risk averse, following Rothschild and Stiglitz (1970). [^6] More formally, we should say that the agent with preferences represented by $\mathcal{U}$ is risk averse, but for simplicity, we say that $\mathcal{U}$ is risk averse. This notion of risk aversion is often called strong risk aversion in the literature, but we omit “strongly” in this paper as we do not deal with the corresponding weak notions. Strict risk aversion holds if $X<_{\rm{cx}}Y$ (meaning $X\leq_{\rm cx}Y$ and $Y\not\leq_{\rm cx}X$) implies $\mathcal{U}(X)>\mathcal{U}(Y)$. Similarly, if a preference functional $\mathcal{U}$ satisfies
$X\leq_{\rm cx}Y~{}\Longrightarrow~{}\mathcal{U}(X)\leq\mathcal{U}(Y),$
then we say that $\mathcal{U}$ is risk seeking, and the strict version is analogous. [^7] It is well known that an EU preference functional with a utility function $u$ is (strictly) risk averse if and only if $u$ is (strictly) concave; it is (strictly) risk seeking if and only if $u$ is (strictly) convex.
Pareto-optimal allocations for risk-averse agents are intrinsically connected to comonotonicity. The two random variables $X,Y$ are said to be comonotonic if
$(X(\omega)-X(\omega^{\prime}))(Y(\omega)-Y(\omega^{\prime}))\geq 0~{}~{}\mbox{for $(\mathbb{P}\times\mathbb{P})$-almost every~{}}(\omega,\omega^{\prime})\in\Omega^{2},$
and a vector of random variables $X_{1},\dots,X_{n}$ is comonotonic if all its component are pairwise comonotonic. Alternatively, the random variables $X_{1},\dots,X_{n}$ are comonotonic if there exists a collection of increasing functions $f_{i}:\mathbb{R}\rightarrow\mathbb{R}$, $i\in[n]$, and a random variable $Z$ such that $X_{i}=f_{i}(Z)$ for all $i\in[n]$ (recall that equalities are in the $\mathbb{P}$-almost sure sense), and we can take $Z=\sum_{i=1}^{n}X_{i}$; see (see Denneberg, 1994, Proposition 4.5).
The comonotonic improvement theorem (in the form of Rüschendorf, 2013, Theorem 10.50) states that, for any $X\in\mathcal{X}=L^{1}$ and any $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, there exists a comonotonic allocation $(Y_{1},\dots,Y_{n})\in\mathbb{A}_{n}(X)$ such that $(Y_{1},\dots,Y_{n})$ is comonotonic and $Y_{i}\leq_{\rm{cx}}X_{i}$ for every $i\in[n]$. As a direct consequence of this result, when all agents are strictly risk averse, Pareto-optimal allocations and equilibrium allocations must be comonotonic. A similar result holds under a different label: in an exchange economy with aggregate risk, the individual consumption of strictly risk-averse EU maximizers is increasing in aggregate wealth.
3.2 Counter-monotonicity and jackpot allocations
Comonotonicity is an extreme type of positive dependence, closely related to risk aversion. When agents are not risk-averse, a form of negative dependence may appear. We first need to define the extreme type of negative dependence.
Two random variables $X,Y$ are counter-monotonic if $X,-Y$ are comonotonic. An allocation $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is (pairwise) counter-monotonic if for every $i\neq j$ the random variables $X_{i},X_{j}$ are counter-monotonic. We omit “pairwise” for simplicity. Unlike comonotonicity, which allows for arbitrary marginal distributions, counter-monotonicity in dimension $n\geq 3$ puts quite strong restrictions on the marginal distributions of the random vector, as shown by Dall’Aglio (1972). Some technical background on counter-monotonicity is collected in Appendix C.
Let $\Pi_{n}$ be the set of all $n$-compositions of $\Omega$, that is,
[…]
In other words, a composition of $\Omega$ is a partition of $\Omega$ with order. [^8] Because we treat almost surely equal random variables as identical, it suffices to require $\mathbb{P}(\bigcup_{i\in[n]}A_{i})=1$ instead of $\bigcup_{i\in[n]}A_{i}=\Omega$. This simplifies some arguments. The indicator function $\mathds{1}_{A}$ for an event $A$ is specified by $\mathds{1}_{A}(\omega)=1$ if $\omega\in A$ and $\mathds{1}_{A}(\omega)=0$ otherwise. Lauzier et al. (2023a, Theorem 1) obtained a stochastic representation of counter-monotonic random vectors $(X_{1},\dots,X_{n})$ with at least three non-constant components, and they have the form $X_{i}=Y\mathds{1}_{A_{i}}+m_{i}$ for some $m_{1},\dots,m_{n}\in\mathbb{R}$, $(A_{1},\dots,A_{n})\in\Pi_{n}$, and $Y\geq 0$ or $Y\leq 0$; a precise statement is reported in Proposition 7 in Appendix C. This reflects the “winner-takes-all” or “loser-loses-all” structure of counter-monotonic allocations. We are particularly interested in the special case
$X_{i}=X\mathds{1}_{A_{i}}~{}\mbox{for all $i\in[n]$, where $(A_{1},\dots,A_{n})\in\Pi_{n}$},$ (2)
where $X=\sum_{i=1}^{n}X_{i}$ and either $X\geq 0$ or $X\leq 0$.
Definition 1.
An allocation $(X_{1},\dots,X_{n})$ is a jackpot allocation if (2) holds for some $X\geq 0$, and it is a scapegoat allocation if (2) holds for some $X\leq 0$.
As we explained above, all counter-monotonic random vectors with at least three non-constant components can be obtained by a constant vector to a jackpot or scapegoat allocation. Therefore, jackpot and scapegoat allocations are representative of extreme negative dependence. Our subsequent study will focus mainly on jackpot allocations for ease of presentation.
A jackpot allocation and a comonotonic allocation are illustrated in Figure 1. Note that in a jackpot allocation, the random vector $(\mathds{1}_{A_{1}},\dots,\mathds{1}_{A_{n}})$ can be correlated with $X$ in an arbitrary way; for instance, it may be independent of $X$ or determined by $X$ (this is the case of Figure 1).
Figure 1: An illustration of a comonotonic allocation $(X/2,X/2)$ of $X$ and a jackpot allocation $(X\mathds{1}_{A},X\mathds{1}_{A^{c}})$ of $X$. In this example, $A$ coincides with the event $\{X\leq 1\}$.
$\omega$payoff001111
(a) $X(\omega)$
$\omega$payoff001111
(b) $X(\omega)/2$
[…]
(c) $(X\mathds{1}_{A})(\omega)$
[…]
(d) $(X\mathds{1}_{A^{c}})(\omega)$
We call $(\mathds{1}_{A_{1}},\dots,\mathds{1}_{A_{n}})$ in (2) a jackpot vector, and denote by $\mathbb{J}_{n}$ the set of all jackpot vectors in $\mathbb{R}^{n}$, that is,
$\mathbb{J}_{n}=\{(\mathds{1}_{A_{1}},\dots,\mathds{1}_{A_{n}}):(A_{1},\dots,A_{n})\in\Pi_{n}\}.$
The set $\mathbb{J}_{n}$ is precisely the set of all random vectors with a generalized Bernoulli distribution (also known as a multinomial distribution with $1$ trial). With this, any jackpot allocation or scapegoat allocation has the form $X\mathbf{J}$ for some $\mathbf{J}\in\mathbb{J}_{n}$. We can check that an equivalent condition for a random vector $(X_{1},\dots,X_{n})$ to be a jackpot allocation is
[…]
Although sharing the formula (2), a jackpot allocation and a scapegoat allocation have very different meanings. In a jackpot allocation, the total wealth $X$ is nonnegative (e.g., a prize), and for each realization of the world $\omega\in\Omega$, only one agent “wins”, i.e., receives all positive payoff, and all other agents receive nothing. In a scapegoat allocation, the total wealth is nonpositive (e.g., a loss), and only one agent “loses”, i.e., suffers the loss. Both types of allocations are often observed in daily life. For instance, the simple lottery ticket (only one winner) is a jackpot allocation, and the “designated driver” of a party is a scapegoat allocation. [^9] By the “designated driver”, we meant the random selection of a participant from the party to be the driver for the evening, who could not drink.
The next result shows a special role of the jackpot and the scapegoat allocations among all counter-monotonic allocations. The probabilistic mixture of two random vectors with joint distributions $F$ and $G$ is another random vector with joint distribution $\lambda F+(1-\lambda)G$ for some $\lambda\in[0,1]$. Probabilistic mixtures are used in the von Neumann–Morgenstern axiomatization of the EU theory, as well as many other models in decision theory (see e.g., Segal, 1990; Wakker, 1994). The next result yields that jackpot allocations are closed under probabilistic mixtures, which directly follows from (3). The same holds for scapegoat allocations by symmetry.
Proposition 1.
A probabilistic mixture of two jackpot allocations is again a jackpot allocation.
In the context of risk sharing, Proposition 1 guarantees that for EU agents, the utility possibility set of all jackpot allocations is a convex set (see Proposition 2 in Section 4). For two general counter-monotonic allocations other than jackpot and scapegoat allocations, their mixture is not necessarily counter-monotonic.
Remark 1.
Our paper concerns decisions under risk. In the context of a decision under ambiguity, there may not be a fixed probability measure $\mathbb{P}$. We can still define a jackpot or scapegoat allocation through (2) or (3), but all equalities need to hold point-wise. In the current definition of jackpot and scapegoat allocations, $\mathbb{P}$ is only used to identify zero-probability events; this is clear from (3).
3.3 Counter-monotonic improvement
A simple assumption of external randomization would be important in the subsequent analysis.
Assumption ER.
There exists a standard uniform (i.e., uniformly distributed on $[0,1]$) random variable $U$ independent of $X$.
Clearly, Assumption ER implies that the probability space $(\Omega,\mathcal{F},\mathbb{P})$ is atomless, meaning that there exists a standard uniform random variable on the space. Intuitively, we can interpret Assumption ER as allowing for an allocation to be implemented with randomization devices like flipping coins, throwing dice, or spinning roulette wheels. The following simple example illustrates the relevance of this assumption.
Example 1.
Suppose that the total wealth is a constant $X=1$, shared by $n$ risk-seeking EU agents, each with a strictly convex utility function. If no external randomization is allowed, then all allocations must be constants, and the allocation $\mathbf{1}/n$ cannot be Pareto-improved for these agents. However, if external randomization is allowed as in Assumption ER, then we can choose $\mathbf{J}\in\mathbb{J}_{n}$ with $\mathbb{E}[\mathbf{J}]=\mathbf{1}/n$, which improves upon $\mathbf{1}/n$, because for any convex function $u$ and event $A$ with $\mathbb{P}(A)=1/n$,
$\mathbb{E}[u(\mathds{1}_{A})]=\frac{1}{n}u(1)+\frac{n-1}{n}u(0)>u(1/n).$
This shows that a randomization by throwing a fair $n$-sided die improves everybody’s utility over the deterministic allocation $\mathbf{1}/n$. In many arguments in the paper, including the proof of Theorem 1 below, the random variable $U$ is used to generate the jackpot vector $\mathbf{J}$.
We are now ready to present our main technical result in this section.
Theorem 1 (Counter-monotonic improvement).
Suppose that $X_{1},\dots,X_{n}\in L^{1}$ are nonnegative, $X=\sum_{i=1}^{n}X_{i}$ and Assumption ER holds. Then, there exists $(Y_{1},\dots,Y_{n})\in\mathbb{A}_{n}(X)$ such that
  1. (i)
    $(Y_{1},\dots,Y_{n})$ is counter-monotonic;
  2. (ii)
    $Y_{i}\geq_{\rm cx}X_{i}$ for $i\in[n]$;
  3. (iii)
    $Y_{1},\dots,Y_{n}$ are nonnegative.
Moreover, $(Y_{1},\dots,Y_{n})$ can be chosen as a jackpot allocation of $X$.
Theorem 1 gives a counterpart to the comonotonic improvement theorem. The importance of Theorem 1 for risk-seeking agents is immediately clear, because $Y_{1},\dots,Y_{n}$ will be preferred by risk-seeking agents over the original payoffs $X_{1},\dots,X_{n}$. Therefore, one may anticipate that for strictly risk-seeking agents, constrained to the set of nonnegative random variables, any Pareto-optimal allocation or equilibrium allocation, if it exists, must be a jackpot allocation; this is formalized in Theorem 2 for EU agents. As another immediate consequence, for any vector of initial endowments, a jackpot allocation obtained in Theorem 1 is individually rational for any risk-seeking agents.
In Theorem 1 and all results later that involve Assumption ER, external randomization is only needed for $X$ but not for $(X_{1},\dots,X_{n})$; that is, the results do not require the following slightly stronger condition:
There exists a standard uniform random variable $U$ independent of $(X_{1},\dots,X_{n})$. (ER*)
Nevertheless, if (ER*) is assumed, then the proof of Theorem 1 can be easily explained. We present a proof of Theorem 1 under this assumption. Write
$Z_{i}=\frac{\sum_{j=1}^{i}X_{j}}{X}\mathds{1}_{\{X>0\}}~{}\mbox{for $i\in[n]$ and $Z_{0}=0$.}$
Define the event $A_{i}=\{Z_{i-1}\leq U<Z_{i}\}$ for $i\in[n]$. Clearly, $A_{1},\dots,A_{n}$ are disjoint and $\mathbb{P}(\bigcup_{i=1}^{n}A_{i})=1$, implying $\sum_{i=1}^{n}\mathds{1}_{A_{i}}=1$ almost surely. Let $Y_{i}=X\mathds{1}_{A_{i}}~{}\mbox{for $i\in[n]$}.$ Clearly, $(Y_{1},\dots,Y_{n})$ is counter-monotonic and it is a jackpot allocation of $X$. For $i\in[n]$, we can derive
$\mathbb{E}\left[Y_{i}\mid X_{1},\dots,X_{n}\right]$ $=\mathbb{E}\left[X\mathds{1}_{\{Z_{i-1}\leq U<Z_{i}\}}\mid X_{1},\dots,X_{n}\right]$
$=\mathbb{E}\left[X(Z_{i}-Z_{i-1})\mid X_{1},\dots,X_{n}\right]$
$=X\frac{X_{i}}{X}\mathds{1}_{\{X>0\}}=X_{i},$
where the last equality follows from the fact that $X_{i}=0$ when $X=0$. Hence, Jensen’s inequality yields $X_{i}\leq_{\rm cx}Y_{i}$. [^10] The conditional expectation relation $X\buildrel\mathrm{d}\over{=}\mathbb{E}[Y|Z]$ for some $Z$ is one of the equivalent conditions for $X\leq_{\rm cx}Y$ used to characterize “increase in risk” by Rothschild and Stiglitz (1970). This proves Theorem 1 under (ER*).
Example 2.
We can apply the above procedure to the constant allocation $(1/n,\dots,1/n)\in\mathbb{A}_{n}(X)$ with $X=1$ in Example 1. Since $X_{i}=1/n$ for $i\in[n]$, we get $Z_{i}=i/n$ for $i\in[n]$. Therefore, the improved allocation $(Y_{1},\dots,Y_{n})$ is given by $Y_{i}=\mathds{1}_{\{(i-1)/n\leq U<i/n\}}$ for $i\in[n]$, which is identically distributed to the allocation $\mathbf{J}$ in Example 1. On the other hand, if we apply the above procedure to a jackpot allocation $(\mathds{1}_{A_{1}},\dots,\mathds{1}_{A_{n}})$, then $Z_{i}=\sum_{j=1}^{i}\mathds{1}_{A_{j}}$, yielding $(Y_{1},\dots,Y_{n})=(\mathds{1}_{A_{1}},\dots,\mathds{1}_{A_{n}})$.
Although the proof of Theorem 2 can be simplified if (ER*) holds, we emphasize that for the risk sharing problems, we cannot make the assumption (ER*), because in that context, the allocation $(X_{1},\dots,X_{n})$ is a decision variable to be chosen by the decision makers. [^11] Technically, this issue can be resolved if we assume that for all random variables, an independent standard uniform random variable exists; however, despite simple intuition, this assumption is very strong and not satisfied by any standard probability space; see Liu et al. (2020, Example 7). Therefore, the difference between (ER*) and Assumption ER is not only a technical concern, but it also affects the applicability of the improvement result in risk sharing.
Assumption ER is not needed for the comonotonic improvement, and the reason is intuitive: for risk-averse agents, external randomization does not enhance their utility, and therefore it is not needed. Mathematically, all comonotonic allocations of $X$ are measurable with respect to the $\sigma$-algebra of $X$ (Denneberg, 1994); this is certainly not true for counter-monotonic allocations.
The assumption that $X_{1},\dots,X_{n}$ are nonnegative is necessary to obtain the existence of jackpot allocations. A similar statement can be made for scapegoat allocations, which then requires $X_{i}\leq 0$ for all $i\in[n]$. The proof follows from observing that in this case, $-X_{1},\dots,-X_{n}$ satisfies the assumptions of Theorem 1. Since convex order is invariant under constant shifts, Theorem 1 and the above statement on the negative risks immediately imply the following result, which has the same form as its comonotonic counterpart.
Corollary 1.
Suppose that $X_{1},\dots,X_{n}\in L^{1}$ are all bounded from above or all bounded from below, $X=\sum_{i=1}^{n}X_{i}$, and Assumption ER holds. Then, there exists a counter-monotonic allocation $(Y_{1},\dots,Y_{n})\in\mathbb{A}_{n}(X)$ such that $Y_{i}\geq_{\rm cx}X_{i}$ for $i\in[n]$.
Whereas results in this section are formulated on $L^{1}$ for generality, in the next few sections we will focus on bounded random variables in the analysis of risk sharing problems.
4 Pareto-optimal allocations for risk-seeking EU agents
We now study Pareto optimality and utility possibility sets for risk-seeking EU agents via the lens of Theorem 1. Competitive equilibria and welfare theorems will be studied in Section 5.
4.1 Setting
We first note that for risk-seeking agents, it is important to set bounds on the set of feasible allocations. Otherwise, a Pareto optimum cannot exist because the two agents will bet on gambles with infinitely large values. For instance, if agents $1$ and $2$ with strictly convexity utility functions $u_{1}$ and $u_{2}$ bet $m$ dollars on an event with $1/2$ probability, then both agents would have infinite expected utility as $m\to\infty$, due to the strict convexity of $u_{1}$ and $u_{2}$. [^12] Note that $u_{1}(m)/2+u_{1}(-m)/2\to\infty$ as $m\to\infty$ if $u_{1}$ is strictly convex. Therefore, any allocation with finite expected utility for agents $1$ and $2$ cannot be Pareto optimal if there are no bounds on the set of allocations.
There are two simple ways to address this issue. The first is to impose lower bounds on the allocation, so $X_{i}\geq a$ for $a\in\mathbb{R}$ and $i\in[n]$. The case $a=0$ is of particular interest because it can be interpreted as a no-short selling/borrowing constraint. The second approach is to restrict the effective domain of $u$; this is a common strategy in the empirical literature, where the power utilities $u(x)=x^{\alpha}$ for $\alpha>0$, are used extensively, and effectively all risks considered are nonnegative. The two approaches are equivalent in our setting, and we formulate them in the following baseline assumption for risk-seeking EU agents.
Assumption EURS.
All agents $i\in[n]$ are EU maximizers with utility function $u_{i}:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ increasing and strictly convex with $u_{i}(0)=0$ for all $i\in[n]$. The domain of allocations is chosen as $\mathcal{X}=L^{\infty}_{+}$, where $L^{\infty}_{+}$ is the set of all nonnegative bounded random variables. The total risk $X\in\mathcal{X}$ satisfies $\mathbb{P}(X>0)>0$.
The assumption that $u_{i}(0)=0$ is a simple normalization, and it does not lose any generality, because a constant shift of the utility function does not change the agent’s preference relation. We also note that since $u_{1},\dots,u_{n}$ are increasing and strictly convex, they are also strictly increasing. The last assumption $\mathbb{P}(X>0)>0$ is made only to exclude the trivial case $X=0$, which only has one possible allocation $(0,\dots,0)$ under the setting $\mathcal{X}=L^{\infty}_{+}$. This will be convenient for us because we will often divide by $\mathbb{E}[u_{i}(X)]>0$.
We will further consider a simpler case of homogeneous risk-seeking EU agents, formally stated below, allowing us to fully characterize the utility possibility set of the risk exchange economy.
Assumption H-EURS.
On top of Assumption EURS, agents are homogeneous, that is, $u_{1}=\dots=u_{n}=u$.
Assumption ER will be used to allow for randomization, which allows us to use Theorem 1. To understand both the Pareto and the competitive equilibria, the function
$V_{{\bm{\lambda}}}:=\max_{i\in[n]}\lambda_{i}u_{i}$
for a given Negishi weight vector ${\bm{\lambda}}=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}_{+}^{n}\setminus\{\mathbf{0}\}$ plays an important role, [^13] The maximum here is taken point-wise, that is, $V_{{\bm{\lambda}}}(x)=\max_{i\in[n]}\lambda_{i}u_{i}(x)$ for all $x\in\mathbb{R}$. and it turns out to be the utility function associated with the representative agent of the $n$ agents. Mathematically, $V_{{\bm{\lambda}}}$ is the upper envelope of functions $\lambda_{1}u_{1},\dots,\lambda_{n}u_{n}$, and it is a convex function.
4.2 Pareto-optimal allocations and the utility possibility frontier
To study Pareto-optimal allocations, we first define the utility possibility set (UPS) and the utility possibility frontier (UPF) of the risk exchange economy. For the total wealth $X$, the UPS is the set of utility vectors $(\mathbb{E}[u_{1}(X_{1})],\dots,\mathbb{E}[u_{n}(X_{n})])$ for $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, denote by $\mathrm{UPS}(X)$. The UPF is the subset of utility vectors achieved by Pareto-optimal allocations, denoted by $\mathrm{UPS}^{\circ}(X)$. As anticipated from Section 3, jackpot allocations play a special role in this economy, and we denote by $\mathrm{UPS}^{\rm J}(X)$ the subset of utility vectors achieved by jackpot allocations. We first establish the convexity of the sets $\mathrm{UPS}^{\rm J}(X)$ and $\mathrm{UPS}(X)$.
Proposition 2.
Under Assumptions ER and EURS, both $\mathrm{UPS}^{\rm J}(X)$ and $\mathrm{UPS}(X)$ are convex.
On the other hand, $\mathrm{UPS}^{\circ}(X)$ is not necessarily convex, and it is the boundary of a convex set. See Example 3 below, where the UPS and UPF are depicted in the case of two agents.
The next result formalizes the observation from the counter-monotonic improvement theorem that Pareto-optimal allocations for strictly risk-seeking agents are jackpot allocations. Moreover, Pareto-optimal allocations are precisely ${\bm{\lambda}}$-optimal allocations.
Theorem 2 (Pareto optimality).
Suppose that Assumptions ER and EURS hold. For an allocation $\mathbf{X}=(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, the following are equivalent:
  1. (i)
    $\mathbf{X}$ is Pareto optimal;
  2. (ii)
    $\mathbf{X}$ is ${\bm{\lambda}}$-optimal for some $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{n})\in\Delta_{n}$;
  3. (iii)
    $\mathbf{X}$ satisfies $\sum_{i=1}^{n}\lambda_{i}\mathbb{E}[u_{i}(X_{i})]=\mathbb{E}\left[V_{\bm{\lambda}}(X)\right]$ for some $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{n})\in\Delta_{n}$,
  4. (iv)
    $\mathbf{X}$ is a jackpot allocation, written as $\mathbf{X}=X(\mathds{1}_{A_{1}},\dots,\mathds{1}_{A_{n}})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}$, such that for some $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{n})\in\Delta_{n}$, $\lambda_{i}u_{i}(X)\mathds{1}_{A_{i}}=V_{{\bm{\lambda}}}(X)\mathds{1}_{A_{i}}$ almost surely for each $i\in[n]$.
We sketch the proof of Theorem 2 by discussing a few key steps. If ${\bm{\lambda}}$ in (ii) has only positive components, then (ii)$\Rightarrow$(i) in Theorem 2 is straightforward; the case with some zero components requires additional arguments. The direction (i)$\Rightarrow$(ii) follows from the Hahn-Banach Theorem and the fact that the utility possibility set is convex by Proposition 2, which relies on Assumption ER. This fact is similar to the convexity of the utility possibility set for concave EU agents (e.g., Mas-Colell et al., 1995 for finite states), but the latter result does not need Assumption ER. Technically, this is because concave utilities get improved by combinations of random variables, but convex utilities do not have such a property and the proof requires a technique of probabilistic mixture via randomization. The equivalence (ii)$\Leftrightarrow$(iii) can be proved by showing that $\mathbb{E}[V_{\bm{\lambda}}(X)]$ is the maximum of $\sum_{i=1}^{n}\lambda_{i}\mathbb{E}[u_{i}(X_{i})]$ over $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, and (i)$\Rightarrow$(iv) can be proved by arguing that the above maximum can only be attained by the jackpot allocations in (iv) with Theorem 1 and some techniques in probability theory.
Theorem 2 immediately yields the relation $\mathrm{UPS}^{\circ}(X)\subseteq\mathrm{UPS}^{\rm J}(X)\subseteq\mathrm{UPS}(X).$ In particular, Pareto-optimal allocations are jackpot allocations. The converse does not hold in general because part (iv) includes a non-trivial condition on $(A_{1},\dots,A_{n})$. The next example illustrates a case in which $\mathrm{UPS}^{\circ}(X)$ is a curve, different from $\mathrm{UPS}^{\rm J}(X)$, which is a convex set.
Example 3.
Set $u_{1}(x)=3x^{2}$ and $u_{2}(x)=4x^{3}$ for $x\geq 0$ and let $X$ be uniformly distributed over $[0,1]$. We have $\mathbb{E}[u_{1}(X)]=\mathbb{E}[u_{2}(X)]=1$, and for any composition $(A_{1},A_{2})$ independent of $X$, it holds that
$\mathbb{E}[u_{1}(X\mathds{1}_{A_{1}})]+\mathbb{E}[u_{2}(X\mathds{1}_{A_{2}})]=\mathbb{P}(A_{1})\mathbb{E}[u_{1}(X)]+\mathbb{P}(A_{2})\mathbb{E}[u_{2}(X)]=1.$
Consider now $A_{1}=\{X\in[0,3/4]\}$ and $A_{2}=\{X\in[3/4,1]\}$ so that $(A_{1},A_{2})$ is a composition that is not independent of $X$. We can compute $\mathbb{E}[u_{1}(X\mathds{1}_{A_{1}})]=\int_{0}^{3/4}3x^{2}\,\mathrm{d}x\approx 0.422$ and $\mathbb{E}[u_{2}(X\mathds{1}_{A_{2}})]=\int_{3/4}^{1}4x^{3}\,\mathrm{d}x\approx 0.684,$ and hence this allocation is better than some jackpot allocations built using events independent of $X$. Intuitively, the allocation that maximizes the equally weighted sum of the welfare gives everything to the agent that has the highest utility pointwise; see the left panel of Figure 2. Maximizing differently weighted sums of the welfare gives the UPF by Theorem 2, plotted in the right panel of Figure 2.
Figure 2: An illustration of Example 3. Left panel: the utility functions. Right panel: the utility possibility frontier.
[…]
0.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111$(0.42,0.68)$$\mathbb{E}[u_{1}(X_{1})]$$\mathbb{E}[u_{2}(X_{2})]$
Although generally $\mathrm{UPS}^{\circ}(X)\neq\mathrm{UPS}^{\rm J}(X)$, there are two special cases in which equality holds and thus all jackpot allocations are Pareto optimal. The first case is described by Assumption H-EURS, where all utility functions are the same, implying that the condition on $(A_{1},\dots,A_{n})$ in part (iv) holds true for ${\bm{\lambda}}=(1/n,\dots,1/n)$ because $V_{{\bm{\lambda}}}(x)=u(x)/n$ for $x\in\mathbb{R}_{+}$. This is formally stated in the result below, which further characterizes the UPF as a simplex.
Theorem 3 (UPF).
Under Assumptions ER and H-EURS, $\mathrm{UPS}^{\circ}(X)=\mathrm{UPS}^{\rm J}(X)=\Delta_{n}(\mathbb{E}[u(X)])$.
Theorem 3 will be useful in Section 5 to unify all four types of allocations: Pareto-optimal, equilibrium, sum-optimal, and jackpot allocations.
The second special case for $\mathrm{UPS}^{\circ}(X)=\mathrm{UPS}^{\rm J}(X)$ is when the aggregate payoff is a constant, as shown in the next proposition, where we do not assume homogeneous agents.
Proposition 3.
If $X=x>0$ is a constant and Assumptions ER and EURS hold, then all jackpot allocations of $X$ are Pareto optimal.
So far in this section, we did not involve the initial endowments, because they are irrelevant for Pareto optimality; yet they are needed for discussing individual rationality. In the setting of this section, individually rational Pareto-optimal allocations always exist (see Lemma 4 in Appendix A), and they are jackpot allocations by Theorem 2.
5 Competitive equilibria and welfare theorems
We now analyze the competitive equilibria of the risk sharing economy, and obtain welfare theorems connecting the Pareto and the competitive equilibria. We continue to work under the setting in Section 4 of risk-seeking EU agents.
5.1 General results
Suppose that our agents satisfy Assumption EURS. Since the set $\mathcal{X}$ of possible risks has only nonnegative elements, the choice for $X_{i}$ needs to satisfy $0\leq X_{i}\leq X$. In other words, no agent can choose a risk $X_{i}$ that is larger than $X$ at any state $\omega\in\Omega$, which is a natural condition. Hence, the individual optimization for agent $i\in[n]$ is
to $\mbox{to maximize}~{}\mathbb{E}[u_{i}(X_{i})]~{}~{}~{}~{}\mbox{over $0\leq X_{i}\leq X$}~{}~{}~{}~{}\mbox{subject to}~{}\mathbb{E}^{Q}[X_{i}]\leq\mathbb{E}^{Q}[\xi_{i}].$ (4)
This corresponds to (1) with $\mathcal{X}_{i}=\{Y:0\leq Y\leq X\}$ and $\mathcal{U}_{i}(Y)=\mathbb{E}[u_{i}(Y)]$ for $Y\in\mathcal{X}$. An equilibrium allocation $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ solves (4) for some $(\xi_{1},\dots,\xi_{n})\in\mathbb{A}_{n}(X)$, which may be left unspecified in the next result.
Theorem 4 (Welfare).
Suppose that Assumptions ER and EURS hold.
  1. (i)
    Every equilibrium allocation of $X$ is Pareto optimal.
  2. (ii)
    Every Pareto-optimal allocation of $X$ is an equilibrium allocation, with an equilibrium price given by
    $\frac{\,\mathrm{d}Q}{\,\mathrm{d}\mathbb{P}}=\frac{V_{\bm{\lambda}}(X)}{X}\frac{1}{\mathbb{E}[V_{\bm{\lambda}}(X)/X]}\mbox{ ~{}~{} with the convention $0/0=0$,}$ (5)
    for some $\bm{\lambda}\in\Delta_{n}$.
By Theorem 2, all Pareto-optimal allocations are jackpot allocations, and hence Theorem 4 (i) yields that all equilibrium allocations are jackpot allocations. The relationship in the two theorems is summarized below:
[…]
$\bm{\lambda}$-optimal allocation (for some ${\bm{\lambda}}$) $\Longleftrightarrow$ Pareto-optimal allocation
The vector ${\bm{\lambda}}$ in (5) is not arbitrarily chosen; indeed, it is the one given in Theorem 2, which will be clear from the proof of Theorem 4. The vector of initial endowments associated with the equilibrium allocation $(X_{1},\dots,X_{n})$ is not unique, and it can be chosen as $(X_{1},\dots,X_{n})$ itself. Another possible choice is the proportional endowments
$\xi_{i}=\frac{\mathbb{E}^{Q}[X_{i}]}{\mathbb{E}^{Q}[X]}X,~{}~{}~{}i\in[n].$
The equilibrium price $Q$ is generally not unique, even for a given vector of initial endowments; the uniqueness issue will be studied in Sections 5.2 and 5.3 below.
The component $\mathbb{E}[V_{\bm{\lambda}}(X)/X]^{-1}$ in (5) is simply a normalization to guarantee that $Q$ is a probability measure. The key property of $Q$ is that, since $u_{1},\dots,u_{n}$ are convex, so is $V_{\bm{\lambda}}$. Hence, $V_{\bm{\lambda}}(x)/x$ is increasing in $x$. This implies that $V_{\bm{\lambda}}(X)/X$ and $X$ are comonotonic. That is, the equilibrium price density $\,\mathrm{d}Q/\,\mathrm{d}\mathbb{P}$ increases as the aggregate endowment becomes more abundant. This comonotonicity property is in stark contrast to the classical setting with strictly risk-averse EU agents, where the equilibrium price is decreasing in the aggregate endowment $X$. Yet, this is not surprising at an intuitive level, as the price structure reflects the marginal utility of agents. For strictly risk-averse EU agents, the marginal utility of wealth is diminishing in total consumption, and so consumption is cheaper in high-endowment states. However, strictly risk-seeking EU agents have a strictly increasing marginal utility, which is reflected in the fact that the unit good is more expensive in high-endowment states.
Example 4.
Consider the case that $u_{i}(x)=a_{i}x$ for $x\geq 0$ and some $a_{i}>0$, $i\in[n]$, so that agents are risk neutral. This is excluded by Assumption EURS (which requires strict convexity), but we can manually check that the statement in (ii) holds true, because equation (5) boils down to $\,\mathrm{d}Q/\,\mathrm{d}\mathbb{P}=1$ and $Q=\mathbb{P}$, which corresponds to the familiar statement that the equilibrium price is the physical probability measure when agents are risk neutral.
The obtained results in this section do not give the existence of a competitive equilibrium for an arbitrary initial endowment vector. While we strongly suspect that any initial allocation $(\xi_{1},\dots,\xi_{n})$ with nonnegative components yields a competitive equilibrium, we did not find a complete proof, and the question remains open. In the next two subsections, we consider two special cases, in which we can explicitly construct the equilibrium.141414 Although we are not able to obtain full characterization on the most general case, a possible approach is discussed in Appendix E, which guarantees that competitive equilibria exist if the initial endowments are proportional to $X$.
5.2 Homogeneous EU agents
We have seen the equivalence between Pareto-optimal allocation, ${\bm{\lambda}}$-optimal allocations, and equilibrium allocations, and all of those allocations are jackpot allocations. For homogeneous agents, the converse direction holds by Theorem 3, as summarized in the next result.
Theorem 5 (Equivalence).
Suppose that Assumptions ER and H-EURS hold. For an allocation $\mathbf{X}=(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, the following are equivalent:
  1. (i)
    $\mathbf{X}$ is a jackpot allocation;
  2. (ii)
    $\mathbf{X}$ is Pareto optimal;
  3. (iii)
    $\mathbf{X}$ is sum-optimal;
  4. (iv)
    $\mathbf{X}$ is an equilibrium allocation;
  5. (v)
    $\sum_{i=1}^{n}\mathbb{E}[u(X_{i})]=\mathbb{E}\left[u(X)\right]$.
Theorem 5 involves a strong structure that is uncommon for EU agents: The equivalence between (ii) and (iii) states that any Pareto-optimal allocation is a maximizer of the social planner’s problem where all the Negishi weights are normalized to one. This equivalence holds also for agents using quantiles or monetary risk measures as their preference functionals; see Embrechts et al. (2018, Proposition 1).
From (v), we see that there is a vector $(\theta_{1},\dots,\theta_{n})\in\Delta_{n}$ such that for every $i\in[n]$ it is $\mathbb{E}[u(X_{i})]=\theta_{i}\mathbb{E}[u(X)]$. The vector $(\theta_{1},\dots,\theta_{n})$ has several interpretations. In the context of (i), consider a jackpot allocation $X\mathbf{J}$ where $\mathbf{J}\in\mathbb{J}_{n}$ is independent of $X$. With the relation $(\theta_{1},\dots,\theta_{n})=\mathbb{E}[\mathbf{J}]$, $\theta_{i}$ can be interpreted as agent $i$’s probability of winning the prize $X$ in a lottery from an independent draw. The special case $\mathbb{E}[\mathbf{J}]=\mathbf{1}/n$ corresponds to a lottery in which every agent has a $1/n$ chance to win. In the context of (iv), for an equilibrium price $Q$, we can let $\theta_{i}=\mathbb{E}^{Q}[\xi_{i}]/\mathbb{E}^{Q}[X]$ and interpret the vector $(\theta_{1},\dots,\theta_{n})$ as a vector of relative purchasing power; this will be made clear in Theorem 6 below.
The relationship in Theorem 5 for homogeneous risk-seeking EU agents is summarized below:
Jackpot allocation $\Longleftrightarrow$ Equilibrium allocation
[…]
Sum-optimal allocation $\Longleftrightarrow$ Pareto-optimal allocation
Another important feature for homogeneous agents is that, for any initial endowment vector $(\xi_{i},\dots,\xi_{n})\in\mathbb{A}_{n}(X)$, we can solve the competitive equilibrium explicitly. Moreover, the equilibrium price $Q$ is unique when at least two agents have non-zero endowments, and it is given by
$\frac{\,\mathrm{d}Q}{\,\mathrm{d}\mathbb{P}}=\frac{u(X)}{X}\frac{1}{\mathbb{E}[u(X)/X]}\mbox{ ~{}~{} with the convention $0/0=0$,}$ (6)
which is precisely (5) in Theorem 4 under Assumption H-EURS. In case only one of $\xi_{1},\dots,\xi_{n}$ is not 00, say $\xi_{1}$, the equilibrium allocation is $(X,0,\dots,0)$, and the equilibrium price is arbitrary.
Theorem 6.
Suppose that Assumptions ER and H-EURS hold. Fix an initial endowment vector $(\xi_{i},\dots,\xi_{n})\in\mathbb{A}_{n}(X)$. Let $Q$ be given by (6), $\theta_{i}=\mathbb{E}^{Q}[\xi_{i}]/\mathbb{E}^{Q}[X]$ for $i\in[n]$, and $\bm{\theta}=(\theta_{1},\dots,\theta_{n}).$
  1. (i)
    The tuple $(X\mathbf{J},Q)$ is a competitive equilibrium for any $\mathbf{J}\in\mathbb{J}_{n}$ independent of $X$ with $\mathbb{E}[\mathbf{J}]=\bm{\theta}$.
  2. (ii)
    The utility vector of any equilibrium allocation $(X_{1},\dots,X_{n})$ is uniquely given by $\mathbb{E}[u(X)]\bm{\theta}$.
  3. (iii)
    If at least two of $\xi_{1},\dots,\xi_{n}$ are not 00, then the equilibrium price $Q$ is uniquely given by (6).
Theorem 6 implies, in particular, that competitive equilibria always exist for any initial endowment vector. Most remarkably, although the equilibrium allocation and its utility vector depend on the initial endowments, the equilibrium price does not. This is different from the case of heterogeneous agents, as (5) depends on the initial endowments implicitly through the parameter ${\bm{\lambda}}\in\Delta_{n}$.
For a given $(\xi_{1},\dots,\xi_{n})$ and a competitive equilibrium $(X\mathbf{J},Q)$, the jackpot vector $\mathbf{J}$ does not need to be independent of $X$ as in part (i) of Theorem 6. Indeed, we can check that any choice of $\mathbf{J}=(J_{1},\dots,J_{n})$ satisfying the budget constraint $\mathbb{E}^{Q}[XJ_{i}]=\mathbb{E}^{Q}[\xi_{i}]$ for $i\in[n]$ is suitable for $(X\mathbf{J},Q)$ to be a competitive equilibrium.
5.3 Two agents
We next obtain that, in the case of any two risk-seeking EU agents, a competitive equilibrium exists for any initial endowments. Moreover, this result also illustrates the non-uniqueness of the equilibrium price.
Proposition 4.
Suppose $n=2$ and Assumptions ER and EURS hold. For any initial endowment vector $(\xi_{1},\xi_{2})\in\mathbb{A}_{2}(X)$, there exists a competitive equilibrium $(X_{1},X_{2},Q)$, where
$\frac{\,\mathrm{d}Q}{\,\mathrm{d}\mathbb{P}}=\frac{u_{1}(X)}{X}\frac{1}{\mathbb{E}[u_{1}(X)/X]}\mbox{ ~{}~{} with the convention $0/0=0$.}$ (7)
The equilibrium price in (7) has the form of (5) with $(\lambda_{1},\lambda_{2})=(1,0)$. Since the positions of agents $1$ and $2$ are symmetric, we immediately get another equilibrium price by replacing $u_{1}$ in (5) with $u_{2}$. Therefore, the equilibrium price is generally not unique, unless $u_{1}=u_{2}$ (seen in Theorem 6). Different equilibrium prices correspond to different equilibrium allocations, but they can be derived from the same vector of initial endowments. Therefore, for a given set of initial endowments, the competitive equilibrium is generally not unique, which is in contrast to the case of strictly risk-averse agents, where often a unique competitive equilibrium can be derived from given initial endowments.
The proof techniques for Proposition 4 do not generalize to the case of $n\geq 3$ agents because our construction of the equilibrium (jackpot) allocation $(X_{1},X_{2})=(X\mathds{1}_{A_{1}},X\mathds{1}_{A_{2}})$ heavily relies on the ratio $u_{1}(x)/u_{2}(x)$. Roughly speaking, we choose $A_{1}$ as the event when $u_{1}(X)/u_{2}(X)$ is large, and choose $A_{2}$ as the event when $u_{1}(X)/u_{2}(X)$ is small; this intuition is similar to part (iv) of Theorem 2, but its generalization to $n\geq 3$ agents is unclear.
6 Rank-dependent utility agents
The previous two sections considered agents that are globally risk seeking. While this assumption was conceptually interesting, more realistic models are needed for applications. We now analyze the problem of sharing risk among RDU agents, which are empirically more relevant. These agents have some local risk-seeking behaviour, but overall they are neither risk seeking nor risk averse. We will mainly focus on Pareto optimality in this section, as deriving competitive equilibrium for RDU agents seems to be beyond the reach of current techniques; an exception is Proposition 5, which offers an explicit competitive equilibria in a special setting. Admittedly, we will make some strong assumptions, but our results lead to interesting new economic insights on risk sharing that were not discovered before. The two main assumptions that we make are: All agents have the same RDU preferences, and their utility function is concave on $\mathbb{R}_{+}$ and linear on a bounded interval $[0,x_{0}]$.
We first describe the RDU model. As in Sections 4 and 5, we consider nonnegative bounded random variables, by letting $\mathcal{X}=L^{\infty}_{+}$. A function $h:[0,1]\rightarrow[0,1]$ is called a probability weighting function if it is increasing and satisfies $h(0)=0$ and $h(1)=1$. For a probability weighting function $h$ and an increasing function $u:\mathbb{R}_{+}\to\mathbb{R}_{+}$ (i.e., a utility function), the RDU preference functional is given by
$\mathcal{U}(Y)=\int u(Y)\,\mathrm{d}(h\circ\mathbb{P})=\int_{0}^{\infty}h(\mathbb{P}(u(Y)>x))\,\mathrm{d}x,~{}~{}~{}Y\in\mathcal{X}.$ (8)
Here, the integral $\int u(Y)\,\mathrm{d}(h\circ\mathbb{P})$ is in the sense of Choquet. An RDU agent has its preference functional given by (8) for some probability weighting function $h$ and utility function $u$. When the utility function $u$ is linear, the RDU agent has the dual utility functional of Yaari (1987), denoted by
$\rho_{h}(Y)=\int Y\,\mathrm{d}(h\circ\mathbb{P}),~{}~{}~{}Y\in\mathcal{X}.$
The RDU functional in (8) is thus given by $\mathcal{U}(Y)=\rho_{h}(u(Y)).$ An RDU agent is risk seeking if $h$ is concave and $u$ is convex, and is risk averse if $h$ is convex and $u$ is concave; see Chew et al. (1987). For nonnegative random variables, the RDU model coincides with the cumulative prospect theory of Tversky and Kahneman (1992) with reference point 00.
The most empirically relevant probability weighting functions $h$ are concave-convex (also called inverse S-shaped), that is, $h$ is continuous and there exists a point $p\in[0,1]$ such that $h$ is concave on $(0,p)$ and convex on $(p,1)$. [^15] If $p=0$ then $h$ is convex, and if $p=1$ then $h$ is concave. Clearly, such a probability weighting function leads to an RDU model that is neither risk averse nor risk seeking. Indeed, the interpretation of such a probability weighting function is risk seeking for small probability events and risk averse for moderate to large probability events, as empirically observed from human behaviour. Tversky and Kahneman (1992) introduced the following class of concave-convex probability weighting functions
$h_{\rm{TK}}(t)=\frac{t^{\gamma}}{(t^{\gamma}+(1-t)^{\gamma})^{1/\gamma}},~{}~{}~{}t\in[0,1],$ (9)
where $\gamma\in(0,1)$ is a parameter, empirically estimated to be $0.71$ by Wu and Gonzalez (1996). [^16] The value of $\gamma$ has been estimated between $0.6$ and $0.8$ from different experiments by many authors including Tversky and Kahneman (1992); see Wu et al. (2004) for a summary.
For a probability weighting function $h$, we denote by $\overline{h}:[0,1]\rightarrow[0,1]$ the concave envelope of $h$, that is,
$\overline{h}(t)=\inf\{g(t):g\geq h\mbox{~{}on [0,1] and $g$ is concave}\},$
which is a concave probability weighting function. Since $\overline{h}\geq h$, we have $\rho_{\overline{h}}(Y)\geq\rho_{h}(Y)$ for all $Y\in\mathcal{X}$. For a concave-convex probability weighting function $h$, its concave envelope $\overline{h}$ coincides with $h$ on an interval $[0,\beta_{h}]\subseteq[0,1]$ (possibly empty) and it is linear on $[\beta_{h},1]$; see panel (b) of Figure 3. With the above background, we now make the main assumption on our homogeneous RDU agents.
Figure 3: An example of a utility function (left panel) and a probability weighting function (right panel) satisfying Assumption H-RDU for $n\geq 8$ and condition (10).
[…]
(a) $u=ax$ on $[0,x_{0}]$ and $u(x)=a\log x+a$ on $[x_{0},\infty)$ for some $a>0$, with $x_{0}=1$
[…]
(b) $h=h_{\rm{TK}}$ with $\gamma=0.71$ (black curve) and $\overline{h}$ (blue curve); $h=\overline{h}$ on $[0,\beta_{h}]$ with $\beta_{h}=0.133$
Assumption H-RDU.
Each agent $i\in[n]$ is an RDU maximizer with utility function $u$ and probability weighting function $h$, where $u$ is increasing and concave on $\mathbb{R}_{+}$ and linear on an interval $[0,x_{0}]$ with $u(0)=0$ and $u(x_{0})>0$, and $h$ is concave-convex with $h=\overline{h}$ for $t\in[0,1/n]$. The domain of allocations is chosen as $\mathcal{X}=L^{\infty}_{+}$. The total risk $X\in\mathcal{X}$ satisfies $\mathbb{P}(X>0)>0$.
Figure 3 depicts one pair $(u,h)$ that satisfies Assumption H-RDU. Using concave $u$ and concave-convex $h$ in RDU models is standard and empirically appealing, as in Tversky and Kahneman (1992). The linearity of $u$ on the interval $[0,x_{0}]$ is technically helpful to employ Theorem 1. The interpretation is $[0,x_{0}]$ represents small-scale payoffs, for which the agent has a linear utility function, which is arguably realistic, as all differentiable utility functions are locally linear. This condition can be generalized to the weaker condition of convexity on $[0,x_{0}]$; see Appendix D. The condition that $h=\overline{h}$ on $[0,1/n]$ is equivalent to $n\geq 1/\beta_{h}$, which is easily satisfied for large $n$. Using $\gamma=0.71$ in (9), we have $h=\overline{h}$ for $t\in[0,0.133]$ as depicted in panel (b) of Figure 3. Here, $n\geq 8$ is sufficient for this condition. This also implies that $h$ is concave on $(0,1/n)$, and we will later use strict concavity of $h$ on $(0,1/n)$ for some statements on strict domination. The conditions on the set $\mathcal{X}$ and the total wealth $X$ are the same as in Assumption EURS.
The next result shows the surprising fact that the previous statement about Pareto optimality for risk-seeking agents can also be obtained for the RDU agents, who are not risk seeking. We will compare a jackpot allocation $X\mathbf{J}$ with the proportional allocation $\overline{\mathbf{X}}:=(X/n,\dots,X/n)$. In the next result, the condition $X\leq x_{0}$ should be interpreted as $\mathbb{P}(X\leq x_{0})=1$, and the following technical condition is used in part (iii):
[…]
In (10), the condition on $h$ holds for $h_{\rm TK}$ in (9), and the condition on $u$ holds for concave utility functions $u$ that are exponential or logarithmic on $[z_{0},\infty)$ for some $z_{0}>0$ (see panel (a) of Figure 3), as well as any bounded utility functions.
Theorem 7 (RDU).
Suppose that Assumptions ER and H-RDU hold. Let $\mathbf{J}\in\mathbb{J}_{n}$ be independent of $X$ satisfying $\mathbb{E}[\mathbf{J}]=\mathbf{1}/n$.
  1. (i)
    If $X\leq x_{0}$, then the jackpot allocation $X\mathbf{J}$ is Pareto optimal and sum-optimal.
  2. (ii)
    If $X\leq x_{0}$, $n\geq 2$, and ${h}$ is strictly concave on $(0,1/n)$, then $X\mathbf{J}$ strictly dominates $\overline{\mathbf{X}}$.
  3. (iii)
    If (10) holds, then there exists $y_{0}>0$ such that for $X\geq y_{0}$, $\overline{\mathbf{X}}$ strictly dominates $X\mathbf{J}$.
  4. (iv)
    If $X$ is a positive constant, $u$ is strictly increasing and differentiable on $\mathbb{R}_{+}$, and ${h}$ is strictly concave on $(0,1/n)$, then $\overline{\mathbf{X}}-\varepsilon\mathbf{1}+n\varepsilon\mathbf{J}$ strictly dominates $\overline{\mathbf{X}}$ for $\varepsilon>0$ small enough.
Part (i) of Theorem 7 gives the Pareto optimality of a jackpot allocation even if the agents are not risk seeking. The key intuition to prove this statement is to notice that, for the jackpot allocation $X\mathbf{J}$, the probability of winning the jackpot is $1/n$, and hence the condition $h=\overline{h}$ on $[0,1/n]$ implies that each RDU agent’s utility for this particular allocation remains the same if we replace $h$ with $\overline{h}$. Then, we can use Theorem 1 to prove that this allocation is indeed optimal for the hypothetical RDU agents using $\overline{h}$, who are risk seeking due to the linear $u$ on $[0,x_{0}]$, and then use this to show the optimality for the non-risk-seeking agents. Parts (iii) and (iv) do not require the assumption on linearity of $u$ on $[0,x_{0}]$, which is clear from the proof.
To interpret Theorem 7, let us consider $X$ as a total prize to share. In parts (i) and (ii), the RDU agents tend to gamble against each other when the total prize is small, i.e., taking values in $[0,x_{0}]$. In particular, such a jackpot allocation is better for every agent when compared with the proportional allocation $\overline{\mathbf{X}}$. However, when the total prize is large enough, as in (iii), it is no longer optimal to use the jackpot allocation, which is strictly dominated by the proportional allocation. It could even be optimal to share proportionally if the utility function has a satiation point (see Example 5 below). Part (iv) considers the case of a constant total wealth $X$ to share. The counter-monotonic (but not jackpot) allocation $\overline{\mathbf{X}}-\varepsilon\mathbf{1}+n\varepsilon\mathbf{J}$ can be interpreted as follows: On top of the constant allocation $\overline{\mathbf{X}}$, every agent pays a fee of $\varepsilon$ to enter a lottery returning $n\varepsilon$ to one randomly selected winner. When $\varepsilon$ is small, this represents a small-stake gamble. Thus, a small-stake gamble is always better than a proportional constant allocation. This observation holds true regardless of the total wealth level, in contrast to parts (i)–(iii).
To summarize these observations, the RDU agents tend to gamble when the stake is small (say, a few hundred dollars), and they may share proportionally, or almost proportionally, when the stake is large (say, several million dollars). This is consistent with empirical observations on risk attitudes: For instance, as observed by Markowitz (1952, p. 153–154), for lottery prizes with equal mean, people tend to choose the risky option for small stakes, and choose the safe option for large stakes. [^17] Markowitz (1952) used a convex-concave utility function in the EU model to explain this for an individual agent. Our utility function is concave, and the gambling behaviour is due to the probability weighting function $h$. If we assume an EU model with a convex-concave utility function of Markowitz (1952) that is convex on $[0,x_{0}]$, then the phenomenon remains to hold; see the proof in Appendix D. A similar empirical evidence is observed by Jullien and Salanié (2000). This result helps to explain the wide existence of small-stake gambles, such as lotteries, sport betting, and casinos.
Example 5.
Assume that $u$ is constant on $[y_{0},\infty)$ for some $y_{0}>x_{0}$; that is, $u$ has a satiation point. If $X\geq ny_{0}$, then $\overline{\mathbf{X}}$ is Pareto optimal because it yields the maximum utility for every agent, and it strictly dominates any jackpot allocation when $h(1/n)<1$ (formally proved in Appendix B).
Some technical discussions are useful. Unlike the case of Theorem 2, we are not able to show that all Pareto-optimal allocations are jackpot allocations or that they are ${\bm{\lambda}}$-optimal; this is due to the non-linearity and nonconvexity in probabilistic mixtures of the RDU functionals. [^18] The functional $\rho_{h}$ is convex in probabilistic mixtures if and only if $h$ is concave; see e.g., Wang et al. (2020, Theorem 3). We do not know whether the UPS is convex. Some extensions of Theorem 7 are discussed in Appendix D. In particular, the conclusions hold for utility functions $u$ such that $x\mapsto u(x)/x$ increases on $[0,x_{0}]$, which is weaker than assuming either linearity or convexity on $[0,x_{0}]$. This allows our result to accommodate convex-concave utility functions, which are found empirically plausible by Jullien and Salanié (2000) using cumulative prospect theory (which is RDU when considering gains).
We have not yet discussed competitive equilibria in this section, because we do not have general results. Nevertheless, in a special case where $X$ is a constant in $(0,x_{0}]$, we can show that $X\mathbf{J}$ is an equilibrium allocation, and the corresponding equilibrium price is the objective probability $\mathbb{P}$.
Proposition 5.
Suppose that Assumptions ER and H-RDU hold, $X=x$ is a constant in $(0,x_{0}]$, and the vector of initial endowments $\bm{\xi}=(\xi_{1},\dots,\xi_{n})\in\mathbb{A}_{n}(x)$ satisfies $\mathbb{E}[\xi_{i}]\leq x\beta_{h}$ for all $i\in[n]$. Then $(x\mathbf{J},\mathbb{P})$ is a competitive equilibrium, for any $\mathbf{J}\in\mathbb{J}_{n}$ satisfying $\mathbb{E}[x\mathbf{J}]=\mathbb{E}[\bm{\xi}]$.
It is not surprising that $\mathbb{P}$ is the equilibrium price, as the total endowment is constant across all states. Although we assume that there is no aggregate uncertainty, the initial endowments in Proposition 5 can be random.
Recall that $\beta_{h}\geq 1/n$ under Assumption H-RDU. Hence, $(\xi_{1},\dots,\xi_{n})=(X/n,\dots,X/n)$ satisfies the condition in Proposition 5, and the corresponding equilibrium allocation is $X\mathbf{J}$ with $\mathbb{E}[\mathbf{J}]=\mathbf{1}/n$ as in Theorem 7. The condition $\mathbb{E}[\xi_{i}]\leq x\beta_{h}$ means that each agent’s initial endowment is not too large compared to the total endowment $x$. Intuitively, agents tend to gamble for small-probability payoffs, which is the case when they have similar initial endowments. On the other hand, if one agent has a relatively large initial endowment, say $0.9x$, then it is no longer optimal for this agent to gamble, because the utility of $0.9x$ is $0.9u(x)$, and the utility of $x\mathds{1}_{A_{1}}$ with $\mathbb{P}(A_{1})=0.9$ is $h(0.9)u(x)$. Typically $h(0.9)<0.9$ (see Figure 3), representing risk aversion for large-probability payoffs. In this case, $(x\mathbf{J},\mathbb{P})$ is not a competitive equilibrium, and we do not know whether equilibria exist.
7 Conclusion
The results in this paper lay a foundation for studying risk sharing with risk-seeking agents and other agents with local risk-seeking behaviour. We anticipate that the counter-monotonic improvement theorem (Theorem 1) will lead to other fruitful directions of future study.
It would be useful to summarize what we know and what we do not. Pareto-optimal allocations and equilibrium allocations for risk-seeking EU agents are fully characterized, and the corresponding welfare theorems are established (Theorems 2 and 4). The case of homogeneous EU agents is even better understood, as we can fully describe the utility possibility frontier (Theorem 3) and the competitive equilibria (Theorems 5 and 6). For a given vector of initial endowment and heterogeneous agents, the existence of competitive equilibria is not proved in general; a potential approach and a special case are discussed in Appendix E. For homogeneous RDU agents with concave-convex probability probability weighting functions, we obtain some Pareto optimal allocations for small-stake payoffs (Theorem 7), but we are not able to offer a complete characterization of Pareto-optimal allocations. We obtain one result on competitive equilibria for RDU agents with no aggregate uncertainty (Proposition 5). General results on competitive equilibria or heterogeneous RDU agents are unavailable. We believe that addressing these open questions requires highly nontrivial analysis.
In most results, we interpreted payoffs and allocations as financial gains. Symmetric results for losses can be derived, with scapegoat allocations in place of jackpot allocations.
To expand the scope, it is natural to investigate whether Pareto-optimal allocations are counter-monotonic in the context of sharing payoffs where agents have heterogeneous beliefs or ambiguity. Intuitively, imposing strong assumptions on the beliefs held by agents can lead to counter-monotonic allocations; there, the agents bet against each other due to belief heterogeneity instead of risk-seeking attitudes. [^19] For instance, in the case of ambiguity-averse agents and no aggregate uncertainty, results in Billot et al. (2000) imply that the emptiness of the intersection of the core of the agents’ capacity is a necessary condition for the Pareto optimality of gambling. We hope that further investigations will shed light on this issue, and we believe that a characterization of the betting behaviour for agents under ambiguity may also benefit from studying jackpot and scapegoat allocations. On this last point, as discussed in Remark 1, the concept of jackpot and scapegoat allocations are well suited for situations with no fixed probability.
Appendices
Appendix A Three technical lemmas
We first provide three technical lemmas that are useful in the proofs of our main results. Denote by $\mathcal{L}$ the set of all random variables $X$ such that there exists a standard uniform random variable $U$ independent of $X$; that is, Assumption ER holds for $X$ being any element of $\mathcal{L}$.
Lemma 1.
Let $X_{1}\in\mathcal{L}$ have distribution $F_{1}$. For any distribution $F$ on $\mathbb{R}^{n}$ with the first marginal $F_{1}$, there exists a random vector $(X_{2},\dots,X_{n})$ such that $(X_{1},X_{2},\dots,X_{n})$ has distribution $F$.
Proof.
Since $X_{1}\in\mathcal{L}$, we can always generate a sequence $U_{2},\dots,U_{n}$ of standard uniform random variables $U_{2},\dots,U_{n}$ from the standard uniform random variable $U_{1}$ independent of $X_{1}$; see e.g., Delbaen (2012, Theorem 1).
Let $(Y_{1},\dots,Y_{n})$ be a random vector with distribution $F$. Let $X_{2}=F^{-1}_{Y_{2}|Y_{1}}(U_{2}|X_{1})$ where $F^{-1}_{Y_{2}|Y_{1}}(\cdot|y)$ is the regular conditional quantile function of $Y_{2}$ given $Y_{1}=y$, which is well defined for almost every $y\in\mathbb{R}$. It is easy to check that $(X_{1},X_{2})\buildrel\mathrm{d}\over{=}(Y_{1},Y_{2})$. Similarly, we can construct $X_{j}$ for $j=3,\dots,n$ by letting
$X_{j}=F_{Y_{j}|Y_{1},\dots,Y_{j-1}}(U_{j}|X_{1},\dots,X_{j-1}),$
where $F_{Y_{j}|Y_{1},\dots,Y_{j-1}}(\cdot|y_{1},\dots,y_{j-1})$ is the quantile function of $Y_{j}$ given $(Y_{1},\dots,Y_{j-1})=(y_{1},\dots,y_{j-1})\in\mathbb{R}^{j-1}$. By construction, we have $(X_{1},\dots,X_{n})\buildrel\mathrm{d}\over{=}(Y_{1},\dots,Y_{n})\sim F$. ∎
Lemma 2.
Fix an atomless probability space. For any $X\in\mathcal{L}$ and $(X_{1},\dots,X_{n})\in\mathcal{X}^{n}$, there exist $(X^{\prime}_{1},\dots,X^{\prime}_{n})\in\mathcal{X}^{n}$ and a standard uniform random variable $U$ such that $(X,X_{1}^{\prime},\dots,X^{\prime}_{n})\buildrel\mathrm{d}\over{=}(X,X_{1},\dots,X_{n})$ and $U$ is independent of $(X,X_{1}^{\prime},\dots,X^{\prime}_{n})$.
Proof.
Let $H:\mathbb{R}^{n+1}\to\mathbb{R}$ be the joint distribution of $(X,X_{1},\dots,X_{n})$. Define $H^{\prime}:\mathbb{R}^{n+2}\to\mathbb{R}$ as $H^{\prime}(x,x_{1},\dots,x_{n},u)=H(x,x_{1},\dots,x_{n})u$. It is clear that $H^{\prime}$ is the joint distribution on $\mathbb{R}^{n+2}$. By Lemma 1, we can find a vector $(X,X^{\prime}_{1},\dots,X^{\prime}_{n},U)\sim H^{\prime}$. Hence, we have $(X,X_{1}^{\prime},\dots,X^{\prime}_{n})\sim H$, $U\sim\mathrm{U}[0,1]$ and $U$ is independent of $(X,X^{\prime}_{1},\dots,X^{\prime}_{n})$. ∎
For the next lemma, recall the definition of $\rho_{h}$, which, for $X\geq 0$, is given by
$\rho_{h}(X)=\int X\,\mathrm{d}(h\circ\mathbb{P})=\int_{0}^{\infty}h(\mathbb{P}(X>x))\,\mathrm{d}x,$
where $h:[0,1]\to[0,1]$ is an increasing function with $h(0)=0=1-h(1)$, called a probability weighting function. For $h$ being the identity on $[0,1]$ we have $\rho_{h}(X)=\mathbb{E}[X].$
Lemma 3.
Suppose that the utility functions $u_{1},\dots,u_{n}$ and the probability weighting functions $h_{1},\dots,h_{n}$ are continuous. For $X\in L_{+}^{\infty}\cap\mathcal{L}$ and $\mathcal{X}=L^{\infty}_{+}$, the set
$\mathrm{UPS}(X)=\Big{\{}\big{(}\rho_{h_{1}}(u_{1}(X_{1})),\dots,\rho_{h_{n}}(u_{n}(X_{2}))\big{)}:(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)\Big{\}}$ (11)
is compact (i.e., closed and bounded). In particular, for $(\lambda_{1},\dots,\lambda_{n})\in\Delta_{n}$, the maximization
$\max_{(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)}\sum_{i=1}^{n}\lambda_{i}\rho_{h_{i}}(u_{i}(X_{i}))$
is attainable.
Proof.
Let $m$ be the essential supremum of $X$. Since distributions in the set
$\{\mbox{distribution of~{}}(X,X_{1},\dots,X_{n}):(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)\}$
are all supported on the bounded region $[0,m]^{n+1}$, it is compact. Hence, any sequence of such distributions has a weak limit, which is supported on the set
$A:=\{(x,x_{1},\dots,x_{n})\in[0,\infty)^{n+1}:x_{1}+\dots+x_{n}=x\}.$
Take any sequence of points $\mathbf{v}_{1},\mathbf{v}_{2},\dots$ in $\mathrm{UPS}(X)$ that converges to $\mathbf{v}_{0}\in\mathbb{R}^{n}$, and let $\mathbf{X}^{(1)},\mathbf{X}^{(2)},\dots$ be the random vectors in $\mathbb{A}_{n}(X)$ with the utility vectors $\mathbf{v}_{1},\mathbf{v}_{2},\dots$, respectively. Let $F$ be a weak limit of the sequence of distributions of the random vectors $(X,\mathbf{X}^{(1)}),(X,\mathbf{X}^{(2)}),\dots$, which we argued above to exist. Note that the first marginal of $F$ is the distribution of $X$. By Lemma 1, there exists a random vector $(X_{1}^{\prime},\dots,X_{n}^{\prime})$ such that $(X,X^{\prime}_{1},\dots,X^{\prime}_{n})$ has distribution $F$. Since $F$ is supported on $A$, we know that $(X^{\prime}_{1},\dots,X^{\prime}_{n})\in\mathbb{A}_{n}(X)$. Moreover, for each $i$, since $u_{i}$ on $[0,\infty)$ and $h_{i}$ on $[0,1]$ are continuous, the function $Y\mapsto\rho_{h_{i}}(u_{i}(Y))$ is continuous with respect to weak convergence by Theorem 4 of Wang et al. (2020). Therefore, $\mathbf{v}_{0}=\lim_{j\to\infty}\mathbf{v}_{j}=(\rho_{h_{1}}(u_{1}(X^{\prime}_{1})),\dots,\rho_{h_{n}}(u_{n}(X^{\prime}_{n})))$, which shows $\mathbf{v}_{0}\in\mathrm{UPS}(X).$ The boundedness statement follows by noting that $\rho_{h_{i}}(u_{i}(X_{i}))\leq\rho_{h_{i}}(u_{i}(X))$ due to monotonicity, which is finite for each $i\in[n]$. The last statement on maximization follows from the compactness of $\mathrm{UPS}(X)$. ∎
Lemma 3 implies in particular that the maxima
[…]
are attainable under Assumptions ER, EURS (for the first max), and H-RDU (for the second max). Moreover, individually rational Pareto-optimal allocations always exist, as summarized in the next lemma.
Lemma 4.
Suppose that Assumption ER hold, and either Assumption EURS or Assumption H-RDU holds. For any initial endowments, individually rational Pareto-optimal allocations exist.
Proof.
Let $(\xi_{1},\dots,\xi_{n})$ be the initial endowment vector. Recall that $\mathcal{U}_{i}$ for $i\in[n]$ is an EU preference functional under Assumption EURS and RDU preference function under Assumption H-RDU. By Lemma 3, the set $\mathrm{UPS}(X)$ is compact. Therefore, the set
$\mathrm{IR}(X):=\{(v_{1},\dots,v_{n})\in\mathrm{UPS}(X):v_{i}\geq\mathcal{U}_{i}(\xi_{i}),~{}i\in[n]\}$
is also compact. Maximizing $\sum_{i=1}^{n}v_{i}$ over $\mathrm{IR}(X)$ yields a point in $\mathrm{UPS}(X)$ attained by an individually rational Pareto-optimal allocation. ∎
Appendix B Proofs of all technical results
B.1 Proofs of results in Section 3
Proof of Proposition 1.
This result follows by noting that the conditions in (3) are preserved under a probabilistic mixture. ∎
Proof of Theorem 1.
We explain that it suffices to prove the result by assuming that there exists a standard uniform random variable $U$ independent of $(X_{1},\dots,X_{n})$. This is due to Lemma 2. To see this, let $(X,X_{1},\dots,X_{n})\buildrel\mathrm{d}\over{=}(X,X_{1}^{\prime},\dots,X_{n}^{\prime})$ be as in Lemma 2. Suppose that we can prove the existence of the desired random vector $(Y_{1},\dots,Y_{n})$ for the vector $(X_{1}^{\prime},\dots,X_{n}^{\prime})$. Note that $\sum_{i=1}^{n}X_{i}^{\prime}=X=\sum_{i=1}^{n}X_{i}$ almost surely. Moreover, $Y_{i}\geq_{\rm cx}X_{i}^{\prime}$ is equivalent to $Y_{i}\geq_{\rm cx}X_{i}$, since $X_{i}\buildrel\mathrm{d}\over{=}X_{i}^{\prime}$. Therefore, $(Y_{1},\dots,Y_{n})$ also satisfies all desired conditions for $(X_{1},\dots,X_{n})$. The rest of the proof follows the arguments in Section 3.3. ∎
Proof of Corollary 1.
Suppose that $X_{1},\dots,X_{n}$ are bounded from below. There exist constants $m_{1},\dots,m_{n}$ such that $X_{1}+m_{1},\dots,X_{n}+m_{n}$ are nonnegative. Write $m=\sum_{i=1}^{n}m_{i}$. By Theorem 1, there exists a jackpot allocation $(Z_{1},\dots,Z_{n})\in\mathbb{A}_{n}(X+m)$ such that $Z_{i}\geq_{\rm cx}X_{i}+m_{i}$ for each $i\in[n]$. It follows that $Z_{i}-m_{i}\geq_{\rm cx}X_{i}$ for $i\in[n]$. Hence, the counter-monotonic random vector $(Z_{1}-m_{1},\dots,Z_{n}-m_{n})$ satisfies the desired conditions in the corollary. The case that $X_{1},\dots,X_{n}$ are bounded from above is analogous. ∎
B.2 Proofs of results in Section 4
Proof of Proposition 2.
We first show the convexity of $\mathrm{UPS}(X)$. Take $\mathbf{x}\in\mathrm{UPS}(X)$ and $\mathbf{y}\in\mathrm{UPS}(X)$. By definition, there exist $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ and $(Y_{1},\dots,Y_{n})\in\mathbb{A}_{n}(X)$ such that
$\mathbf{x}=\left(\mathbb{E}\left[u_{1}(X_{1})\right],\dots,\mathbb{E}\left[u_{n}(X_{n})\right]\right)\mbox{ and }\mathbf{y}=\left(\mathbb{E}\left[u_{1}(Y_{1})\right],\dots,\mathbb{E}\left[u_{n}(Y_{n})\right]\right).$
By Lemma 2, we can assume, without loss of generality, that there exists a uniform random variable $U$ independent of $(X,X_{1},\dots,X_{n})$ and $(X,Y_{1},\dots,Y_{n})$; otherwise we can use $(X,X^{\prime}_{1},\dots,X^{\prime}_{n})$ and $(X,Y^{\prime}_{1},\dots,Y^{\prime}_{n})$ as in Lemma 2, and they will yield the same vectors of utility as $\mathbf{x}$ and $\mathbf{y}$.
Consider their convex combination $\alpha\mathbf{x}+(1-\alpha)\mathbf{y}$ for some $\alpha\in[0,1]$. Let $Z_{i}=X_{i}\mathds{1}_{\{U\leq\alpha\}}+Y_{i}\mathds{1}_{\{U>\alpha\}}$ for $i\in[n]$. Clearly, $\sum_{i=1}^{n}Z_{i}=X$; that is, $(Z_{1},\dots,Z_{n})\in\mathbb{A}_{n}(X)$. Moreover, for every $i\in[n]$, we have
$\mathbb{E}[u_{i}(Z_{i})]$ $=\mathbb{E}\left[u_{i}\left(X_{i}\mathds{1}_{\{U\leq\alpha\}}+Y_{i}\mathds{1}_{\{U>\alpha\}}\right)\right]$
$=\mathbb{E}\left[u_{i}(X_{i})\mathds{1}_{\{U\leq\alpha\}}+u_{i}(Y_{i})\mathds{1}_{\{U>\alpha\}}\right]=\alpha\mathbb{E}[u_{i}(X_{i})]+(1-\alpha)\mathbb{E}[u_{i}(Y_{i})],$
Thus, $\alpha\mathbf{x}+(1-\alpha)\mathbf{y}\in\mathrm{UPS}(X)$, and $\mathrm{UPS}(X)$ is a convex set.
To show the convexity of $\mathrm{UPS}^{\rm J}(X)$, we follow the same arguments as above, with the additional observation that by using Proposition 1 on mixtures of jackpot allocations, $(Z_{1},\dots,Z_{n})$ constructed above is a jackpot allocation if $(X_{1},\dots,X_{n})$ and $(Y_{1},\dots,Y_{n})$ are both jackpot allocations. ∎
Before proving Theorem 2, we first present a lemma, which gives (ii)$\Leftrightarrow$(iii) in Theorem 2.
Lemma 5.
Suppose that Assumption EURS hold and ${\bm{\lambda}}=(\lambda_{1},\dots,\lambda_{n})\in\Delta_{n}$. Then,
$\max_{(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)}\sum_{i=1}^{n}\lambda_{i}\mathbb{E}[u_{i}(X_{i})]=\mathbb{E}\left[V_{\bm{\lambda}}(X)\right].$
Proof of Lemma 5.
By Lemma 3, $\max_{(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)}\sum_{i=1}^{n}\lambda_{i}\mathbb{E}[u_{i}(X_{i})]$ is attainable. Take any composition $(A_{1},\dots,A_{n})\in\Pi_{n}$ satisfying $A_{i}\subseteq\{\lambda_{i}u_{i}(X)=V_{\bm{\lambda}}(X)\}$ for $i\in[n]$. An explicit choice is
$A_{i}=\left\{\omega\in\Omega:i=\min\left(\operatorname*{arg\,max}_{j\in[n]}\lambda_{j}u_{j}\left(X(\omega)\right)\right)\right\},~{}~{}i\in[n].$
By construction, we have
$\max_{(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)}\sum_{i=1}^{n}\lambda_{i}\mathbb{E}[u_{i}(X_{i})]$ $\geq\sum_{i=1}^{n}\mathbb{E}\left[\lambda_{i}u_{i}\left(X\mathds{1}_{A_{i}}\right)\right]$
$=\sum_{i=1}^{n}\mathbb{E}\left[\lambda_{i}u_{i}(X)\mathds{1}_{A_{i}}\right]=\sum_{i=1}^{n}\mathbb{E}\left[V_{\bm{\lambda}}(X)\mathds{1}_{A_{i}}\right]=\mathbb{E}\left[V_{\bm{\lambda}}(X)\right].$
Moreover, for any allocation $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, we have
$\sum_{i=1}^{n}\mathbb{E}[\lambda_{i}u_{i}(X_{i})]\leq\sum_{i=1}^{n}\mathbb{E}[V_{\bm{\lambda}}(X_{i})]\leq\mathbb{E}\left[V_{\bm{\lambda}}\left(\sum_{i=1}^{n}X_{i}\right)\right]=\mathbb{E}[V_{\bm{\lambda}}(X)],$
where the last inequality follows from the superadditivity of $V_{\bm{\lambda}}$ because $V_{\bm{\lambda}}$ is convex with $V_{\bm{\lambda}}(0)=0$. Combining the above two inequalities, we get the desired result. ∎
Proof of Theorem 2.
We prove the theorem following the route (i)$\Rightarrow$(ii)$\Rightarrow$(iii)$\Rightarrow$(i)$\Rightarrow$(iv)$\Rightarrow$(iii).
(i)$\Rightarrow$(ii): By Proposition 2, the UPS of $X$ is a convex set. By the Hahn-Banach Theorem, for every Pareto-optimal allocation $\mathbf{X}$, there exists $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{n})\in\Delta_{n}$ such that $\mathbf{X}$ is a ${\bm{\lambda}}$-optimal allocation.
The equivalence (ii)$\Leftrightarrow$(iii) follows from Lemma 5.
(iii)$\Rightarrow$(i): If ${\bm{\lambda}}$ has only positive components, it is clear that ${\bm{\lambda}}$-optimality implies Pareto optimality. Now suppose that some components of ${\bm{\lambda}}$ are 00. Without loss of generality, suppose that $\lambda_{1},\dots,\lambda_{j}$ are positive, and $\lambda_{j+1},\dots,\lambda_{n}$ are zero. Let $\mathbf{X}=(X_{1},\dots,X_{n})$ satisfy $\sum_{i=1}^{n}\lambda_{i}\mathbb{E}[u_{i}(X_{i})]=\mathbb{E}[V_{\bm{\lambda}}(X)]$ and suppose that $\mathbf{Y}=(Y_{1},\dots,Y_{n})$ strictly dominates $\mathbf{X}$. Note that ${\bm{\lambda}}$-optimality of $\mathbf{X}$ implies that $\mathbb{E}[u_{i}(X_{i})]=\mathbb{E}[u_{i}(Y_{i})]$ for $i\in[j]$. Hence, $\mathbb{E}[u_{i}(Y_{i})]>\mathbb{E}[u_{i}(X_{i})]$ must hold for some $i\geq j+1$, and this further implies $\mathbb{P}(\sum_{i=1}^{j}Y_{i}<X)>0$. Write $Y=\sum_{i=1}^{j}Y_{i}$. Note that $V_{\bm{\lambda}}$ is the maximum of $j$ strictly increasing functions, and hence it is strictly increasing. Using $Y\leq X$ and $\mathbb{P}(Y<X)>0$, we have $\mathbb{E}[V_{{\bm{\lambda}}}(Y)]<\mathbb{E}[V_{{\bm{\lambda}}}(X)]$. By Lemma 5 and $\lambda_{i}=0$ for $i\geq j+1$, we have
$\sum_{i=1}^{j}\mathbb{E}[\lambda_{i}u_{i}(Y_{i})]\leq\mathbb{E}\left[V_{{\bm{\lambda}}}(Y)\right]<\mathbb{E}\left[V_{{\bm{\lambda}}}(X)\right]=\sum_{i=1}^{j}\mathbb{E}[\lambda_{i}u_{i}(X_{i})],$
contradicting that $\mathbf{Y}$ dominates $\mathbf{X}$. Hence, $\mathbf{X}$ is Pareto optimal.
(i)$\Rightarrow$(iv): By Theorem 1, there is a jackpot allocation $\mathbf{Y}=(Y_{1},\dots,Y_{n})\in\mathbb{A}_{n}(X)$ such that $Y_{i}\geq_{\rm cx}X_{i}$ holds for all $i\in[n]$. As $u_{i}$ is strictly convex, we have $\mathbb{E}[u(Y_{i})]=\mathbb{E}[u(X_{i})]$ by Pareto optimality of $(X_{1},\dots,X_{n})$. By Shaked and Shanthikumar (2007, Theorem 3.A.43), we obtain $Y_{i}\buildrel\mathrm{d}\over{=}X_{i}$. Next we show $\mathbf{X}\buildrel\mathrm{d}\over{=}\mathbf{Y}$. By Puccetti and Wang (2015, Theorem 3.3), $(Y_{1},\dots,Y_{n})$ solves
$\inf\left\{\mathbb{E}[(Z_{1}+\dots+Z_{n})^{2}]:~{}Z_{i}\buildrel\mathrm{d}\over{=}Y_{i},~{}i\in[n]\right\}.$ (12)
Since $X_{i}\buildrel\mathrm{d}\over{=}Y_{i}$, $i\in[n]$ and $\sum_{i=1}^{n}X_{i}=\sum_{i=1}^{n}Y_{i}$, it is clear that $(X_{1},\dots,X_{n})$ also solves (12). Since $\mathcal{X}=L^{\infty}$, $\mathbb{E}[|Y_{i}|]<\infty$ for $i\in[n]$. Hence, we have that $(X_{1},\dots,X_{n})$ is counter-monotonic by Puccetti and Wang (2015, Theorem 3.3), which further gives $\mathbf{X}\buildrel\mathrm{d}\over{=}\mathbf{Y}$. Write
$(X_{1},\dots,X_{n})=(X\mathds{1}_{A_{1}},\dots,X\mathds{1}_{A_{n}})~{}~{}\mbox{for some}~{}~{}(A_{1},\dots,A_{n})\in\Pi_{n}.$
Using (iii), which is implied by (i), $\sum_{i=1}^{n}\lambda_{i}\mathbb{E}[u_{i}(X_{i})]=\mathbb{E}[V_{\bm{\lambda}}(X)]$ for some ${\bm{\lambda}}=(\lambda_{1},\dots,\lambda_{n})\in\Delta_{n}$. It is clear that $\lambda_{i}u_{i}(X_{i})=\lambda_{i}u_{i}(X)\mathds{1}_{A_{i}}\leq V_{\bm{\lambda}}(X)\mathds{1}_{A_{i}}$ for all $i\in[n]$. If there exists $i\in[n]$ such that $\mathbb{P}(\lambda_{i}u_{i}(X)\mathds{1}_{A_{i}}<V_{\bm{\lambda}}(X)\mathds{1}_{A_{i}})>0$, then $\mathbb{E}[\lambda_{i}u_{i}(X)\mathds{1}_{A_{i}}]<\mathbb{E}[V_{\bm{\lambda}}(X)\mathds{1}_{A_{i}}]$. Hence,
$\sum_{i=1}^{n}\lambda_{i}\mathbb{E}[u_{i}(X_{i})]<\sum_{i=1}^{n}\lambda_{i}\mathbb{E}[V_{\bm{\lambda}}(X)\mathds{1}_{A_{i}}]=\mathbb{E}\left[V_{\bm{\lambda}}(X)\right],$
contradicting ${\bm{\lambda}}$-optimality. Hence, (iv) holds.
The remaining (iv)$\Rightarrow$(iii) follows by noting $\sum_{i=1}^{n}\lambda_{i}u_{i}(X_{i})=\sum_{i=1}^{n}\lambda_{i}V_{{\bm{\lambda}}}(X)\mathds{1}_{A_{i}}=V_{{\bm{\lambda}}}(X).$
Proof of Theorem 3.
We first show that $\mathrm{UPS}^{\rm J}(X)\subseteq\Delta_{n}(\mathbb{E}[u(X)])$. Let $(X_{1},\dots,X_{n})$ be a jackpot allocation of $X$ and observe that by construction we have $\mathbb{E}[u(X_{i})]\geq 0$ for $i\in[n]$, and
$\sum_{i=1}^{n}\mathbb{E}[u(X_{i})]=\sum_{i=1}^{n}\mathbb{E}[u(X)\mathds{1}_{A_{i}}]=\mathbb{E}[u(X)].$
Hence, $(\mathbb{E}[u(X_{1})],\dots,\mathbb{E}[u(X_{n})])\in\Delta_{n}(\mathbb{E}[u(X)])$. Conversely, let $(\theta_{1},\dots,\theta_{n})\in\Delta_{n}$. By Assumption ER, we can take $(A_{1},\dots,A_{n})\in\Pi_{n}$ independent of $X$ such that $\mathbb{P}(A_{i})=\theta_{i}$ for $i\in[n]$. This gives $\mathbb{E}[u(X\mathds{1}_{A_{i}})]=\mathbb{P}(A_{i})\mathbb{E}[u(X)]=\theta_{i}\mathbb{E}[u(X)]$ for $i\in[n]$. Therefore, $\mathrm{UPS}^{\rm J}(X)=\Delta_{n}(\mathbb{E}[u(X)])$. Note that every point in $\Delta_{n}(\mathbb{E}[u(X)])$ is in the UPF as they are not dominated by any other points in $\mathrm{UPS}^{\rm J}(X)$. Together with $\mathrm{UPS}^{\circ}(X)\subseteq\mathrm{UPS}^{\rm J}(X)$ guaranteed by Theorem 2, we can conclude that $\mathrm{UPS}^{\circ}(X)=\mathrm{UPS}^{\rm J}(X)=\Delta_{n}(\mathbb{E}[u(X)])$. ∎
Proof of Proposition 3.
Suppose that a jackpot allocation $x\mathbf{J}$ for $\mathbf{J}\in\mathbb{J}_{n}$ is strictly dominated by an allocation $\mathbf{Y}$. By Theorem 1, $\mathbf{Y}$ is dominated by another jackpot allocation $x\mathbf{J}^{\prime}$ with $\mathbf{J}^{\prime}\in\mathbb{J}_{n}$. The strict domination of $x\mathbf{J}^{\prime}$ over $x\mathbf{J}$ implies $\mathbb{E}[x\mathbf{J}^{\prime}]\geq\mathbb{E}[x\mathbf{J}]$ componentwise with at least one component having strict inequality. This is not possible because both $\mathbf{J}$ and $\mathbf{J}^{\prime}$ have components summing to $1$. Hence, $x\mathbf{J}$ cannot be strictly dominated by $\mathbf{Y}$, and it is Pareto optimal. ∎
B.3 Proofs of results in Section 5
Proof of Theorem 4.
(i) Suppose that $(X_{1},\dots,X_{n},Q)$ is an equilibrium but $(X_{1},\dots,X_{n})$ is strictly dominated by another allocation $(Y_{1},\dots,Y_{n})\in\mathbb{A}_{n}(X)$. There exists $j\in[n]$ such that $\mathbb{E}[u_{j}(Y_{j})]>\mathbb{E}[u_{j}(X_{j})]$, and by the fact that $(X_{1},\dots,X_{n},Q)$ is an equilibrium, it must hold that $\mathbb{E}^{Q}[Y_{j}]>\mathbb{E}^{Q}[\xi_{i}]\geq\mathbb{E}^{Q}[X_{j}]$. Since $\sum_{i=1}^{n}\mathbb{E}^{Q}[Y_{i}]=\mathbb{E}^{Q}[X]=\sum_{i=1}^{n}\mathbb{E}^{Q}[X_{i}],$ there exists $i\in[n]$, $i\neq j$ such that $\mathbb{E}^{Q}[Y_{i}]<\mathbb{E}^{Q}[X_{i}]$. By Pareto dominance we also have $\mathbb{E}[u_{i}(Y_{i})]\geq\mathbb{E}[u_{i}(X_{i})]$. Let $\alpha=\mathbb{E}[X_{i}-Y_{i}]/\mathbb{E}[X-Y_{i}]$. Since $\mathbb{E}^{Q}[Y_{i}]<\mathbb{E}^{Q}[X_{i}]\leq\mathbb{E}^{Q}[X]$, we have $\alpha\in(0,1]$. Let
$Z_{i}=Y_{i}+(X-Y_{i})\alpha.$
It is clear that $Y_{i}\leq Z_{i}\leq X$ and $\mathbb{E}^{Q}[Z_{i}]=\mathbb{E}^{Q}[X_{i}]\leq\mathbb{E}^{Q}[\xi_{i}]$. Recall that $\mathbb{E}^{Q}[Y_{i}]<\mathbb{E}^{Q}[X]$, which implies $Q(Z_{i}>Y_{i})>0$, and hence, $\mathbb{P}(Z_{i}>Y_{i})>0$. Since $u_{i}$ is strictly increasing, we obtain $\mathbb{E}[u_{i}(Z_{i})]>\mathbb{E}[u_{i}(Y_{i})]\geq\mathbb{E}[u_{i}(X_{i})]$, contradicting individual optimality for agent $i$.
(ii) Let $\mathbf{X}=(X_{1},\dots,X_{n})$ be a Pareto-optimal allocation. By Theorem 2, we have $\mathbf{X}=X\mathbf{J}$ for some $\mathbf{J}=(\mathds{1}_{A_{1}},\dots,\mathds{1}_{A_{n}})\in\mathbb{J}_{n}$ and $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{n})\in\Delta_{n}$ such that $(\lambda_{1}u_{1}(X_{1}),\dots,\lambda_{n}u_{n}(X_{n}))=V_{\bm{\lambda}}(X)\mathbf{J}$. From the proof of Theorem 2, we can take $\lambda_{i}=0$ if $X_{i}=0$ (almost surely, omitted) and $\lambda_{i}>0$ if $X_{i}$ is not 0.
Take the initial endowment $(\xi_{1},\dots,\xi_{n})=(X_{1},\dots,X_{n})$. We will show that $(X_{1},\dots,X_{n},Q)$ is a competitive equilibrium for $(\xi_{1},\dots,\xi_{n})$. Let $z=\mathbb{E}[V_{\bm{\lambda}}(X)/X]$ and $Q$ be given by (5). If $\xi_{i}=0$, it is clear that $X_{i}=0$ solves individual optimality. Next, we discuss the case that $\xi_{i}$ is not 0, which implies $\lambda_{i}>0$. For such $i\in[n]$ let
$x_{i}=\mathbb{E}^{Q}[\xi_{i}]=\mathbb{E}^{Q}[X_{i}]=\mathbb{E}^{Q}[X\mathds{1}_{A_{i}}],$
and notice that
$\lambda_{i}\mathbb{E}[u_{i}(X_{i})]=\mathbb{E}\left[V_{\bm{\lambda}}(X)\mathds{1}_{A_{i}}\right]=\mathbb{E}\left[X\frac{V_{\bm{\lambda}}(X)}{X}\mathds{1}_{A_{i}}\right]=z\mathbb{E}^{Q}\left[{X}\mathds{1}_{A_{i}}\right]=zx_{i}.$
For any $Y_{i}$ satisfying $0\leq Y_{i}\leq X$ and the budget constraint $\mathbb{E}^{Q}\left[Y_{i}\right]\leq{x_{i}}$ we have
$\lambda_{i}\mathbb{E}[u_{i}(Y_{i})]$ $=\mathbb{E}\left[Y_{i}\frac{\lambda_{i}u_{i}(Y_{i})}{Y_{i}}\right]\leq\mathbb{E}\left[Y_{i}\frac{\lambda_{i}u_{i}(X)}{X}\right]\leq\mathbb{E}\left[Y_{i}\frac{V_{\bm{\lambda}}(X)}{X}\right]=x\mathbb{E}^{Q}[Y_{i}]\leq zx_{i}=\lambda_{i}\mathbb{E}[u(X_{i})],$
where the first inequality using the fact that $x\mapsto u(x)/x$ is increasing and the second that $\lambda_{i}u_{i}\leq V_{\bm{\lambda}}$. Therefore, $(X_{1},\dots,X_{n},Q)$ satisfies individual optimality. Market clearance also holds, because $\sum_{i=1}^{n}\mathds{1}_{A_{i}}=1.$ Therefore, $(X_{1},\dots,X_{n})$ is an equilibrium allocation with itself being the initial endowments. ∎
Proof of Theorem 5.
The equivalence between (i), (ii), (iii) and (v) follows from Theorems 2 and 3. The equivalence between them and (iv) follows from Theorem 4. ∎
Proof of Theorem 6.
We first show (i). Denote by $x_{i}=\mathbb{E}^{Q}[\xi_{i}]=\theta_{i}\mathbb{E}^{Q}[X]$ and $z=\mathbb{E}[u(X)/X]>0$. It follows that
$\mathbb{E}^{Q}[X_{i}]=\mathbb{E}^{Q}[X\mathds{1}_{A_{i}}]=\mathbb{E}[u(X)\mathds{1}_{A_{i}}]=\mathbb{E}^{Q}[X]\mathbb{P}(A_{i})=x_{i},$
and hence, the budget constraint is satisfied for each $i\in[n]$. Moreover,
$\mathbb{E}[u(X_{i})]=\mathbb{E}[u(X)\mathds{1}_{A_{i}}]=\mathbb{E}\left[X\frac{u(X)}{X}\mathds{1}_{A_{i}}\right]=z\mathbb{E}^{Q}[X\mathds{1}_{A_{i}}]=zx_{i}.$
For any $Y_{i}$ satisfying $0\leq Y_{i}\leq X$ and the budget constraint $\mathbb{E}^{Q}\left[Y_{i}\right]\leq{x_{i}}$, using the fact that $x\mapsto u(x)/x$ is increasing, we have
$\mathbb{E}[u(Y_{i})]=\mathbb{E}\left[Y_{i}\frac{u(Y_{i})}{Y_{i}}\right]\leq\mathbb{E}\left[Y_{i}\frac{u(X)}{X}\right]=z{\mathbb{E}^{Q}\left[Y_{i}\right]}\leq z{x_{i}}=\mathbb{E}[u(X_{i})].$
Therefore, $(X_{1},\dots,X_{n},Q)$ satisfies individual optimality. Market clearance also holds, because $\sum_{i=1}^{n}\mathds{1}_{A_{i}}=1.$
Next, we show part (iii), the uniqueness of the equilibrium price. Suppose that $(X_{1},\dots,X_{n},Q)$ is a competitive equilibrium. By Theorem 5 we know it is a jackpot allocation. We can write $(X_{1},\dots,X_{n})=(X\mathds{1}_{A_{1}},\dots,X\mathds{1}_{A_{n}})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}$ satisfying $\mathbb{E}^{Q}[X_{i}]=\mathbb{E}^{Q}[\xi_{i}]$ by the binding budget constraint. Let $P$ be the conditional probability measure of $\mathbb{P}$ on $\{X>0\}$, and let $p=\mathbb{P}(X>0)>0$. Let $\eta=\,\mathrm{d}Q/\,\mathrm{d}P$ and define a probability measure $R$ by
$\frac{\,\mathrm{d}R}{\,\mathrm{d}Q}=\frac{X}{c},~{}\mbox{where $c=\mathbb{E}^{Q}[X]$},$
and note that for any $A\in\mathcal{F}$, we have
$\mathbb{E}[u(X\mathds{1}_{A})]=\frac{1}{p}\mathbb{E}^{P}[u(X\mathds{1}_{A})]=\frac{1}{p}\mathbb{E}^{R}\left[\frac{\,\mathrm{d}P}{\,\mathrm{d}Q}\frac{\,\mathrm{d}Q}{\,\mathrm{d}R}u(X)\mathds{1}_{A}\right]=\frac{1}{p}\mathbb{E}^{R}\left[\frac{cu(X)}{\eta X}\mathds{1}_{A}\right].$
Denote by $Z=cu(X)/(\eta X)$. Individual optimality of $(X_{1},\dots,X_{n})$ implies that for any $i\in[n]$ and any $A\in\mathcal{F}$ satisfying $\mathbb{E}^{Q}[X\mathds{1}_{A}]\leq\mathbb{E}^{Q}[X\mathds{1}_{A_{i}}]$, we have
$\mathbb{E}^{R}\left[Z\mathds{1}_{A}\right]=p\mathbb{E}[u(X\mathds{1}_{A})]\leq p\mathbb{E}[u(X\mathds{1}_{A_{i}})]=\mathbb{E}^{R}\left[Z\mathds{1}_{A_{i}}\right].$
Note that $\mathbb{E}^{Q}[X\mathds{1}_{A}]\leq\mathbb{E}^{Q}[X\mathds{1}_{A_{i}}]$ is equivalent to $R(A)\leq R(\mathds{1}_{A_{i}})$. Take $A\in\mathcal{F}$ such that $R(A)=R({A_{i}})$ and $Z$ and $\mathds{1}_{A}$ are comonotonic; this is possible because $R$ is atomless. Suppose for contradiction that $Z$ is not a constant. The Fréchet-Hoeffding inequality gives
$\mathrm{cov}(Z,\mathds{1}_{A_{i}})\geq\mathrm{cov}(Z,\mathds{1}_{A})\geq 0,$
and $\mathrm{cov}(Z,\mathds{1}_{A})>0$ if $R({A})\in(0,1)$. Since at least two of $\xi_{1},\dots,\xi_{n}$ are not 00, by the binding budget constraint, at least two of ${A_{1}},\dots,{A_{n}}$ have positive probability under $R$. Therefore, $\mathrm{cov}(Z,\mathds{1}_{A_{i}})>0$ for at least one $i$. However, $\sum_{i=1}^{n}\mathrm{cov}(Z,\mathds{1}_{A_{i}})=\mathrm{cov}(Z,1)=0$, a contradiction. Hence, $Z$ is a constant. Therefore, $\eta$ is equal to a constant times $u(X)/X$, showing that $Q$ has the form (6).
Finally, part (ii) follows immediately from (iii), by noting that for the unique price $Q$, there can only be one maximum utility value for every agent. If only one of $\xi_{1},\dots,\xi_{n}$ is not zero, say $\xi_{1}=X$, then the allocation is uniquely given by $(X,0,\dots,0)$, and hence the unique vector of utilities. ∎
Proof of Proposition 4.
Without loss of generality, we can assume $\mathbb{E}[\xi_{1}]>0$ and $\mathbb{E}[\xi_{2}]>0$; otherwise $(X,0)$ or $(0,X)$ is an equilibrium allocation with any equilibrium price absolutely continuous with respect to $\mathbb{P}$. Moreover, we can assume $\mathbb{P}(X=0)=0$, because the allocation on the event $\{X=0\}$ is trivial.
For a random variable $W$, a tail event is an event $A$ such that for some $w\in\mathbb{R}$, $W\geq w$ on $A$ and $W\leq w$ on $A^{c}$. In an atomless probability space, a tail event with any given probability $\lambda\in(0,1)$ exists, as shown by Wang and Zitikis (2021). Let $W=u_{1}(X)/u_{2}(X)$. For $\lambda\in(0,1)$, let $(A^{\lambda})_{\lambda\in(0,1)}$ be an increasing family of tail events of $W$ such that $\mathbb{P}(A^{\lambda})=\lambda$. We can check that the mapping $\lambda\mapsto\mathbb{E}^{Q}[X\mathds{1}_{A^{\lambda}}]$ is continuous (because $\lambda\mapsto Q(A^{\lambda})$ is continuous) and its range is the open interval $(0,\mathbb{E}^{Q}[X])$. Therefore, there exists $\lambda^{*}\in(0,1)$ such that $\mathbb{E}^{Q}[X\mathds{1}_{A^{\lambda}}]=\mathbb{E}^{Q}[\xi_{1}]\in(0,\mathbb{E}^{Q}[X])$. Write $A_{1}=A^{\lambda}$ and $A_{2}=(A^{\lambda})^{c}$. By definition of the tail event, for some $w^{*}\geq 0$, we have $W\geq w^{*}$ on $A_{1}$ and $W\leq w^{*}$ on $A_{2}$.
We will show that $(X_{1},X_{2})=(X\mathds{1}_{A_{1}},X\mathds{1}_{A_{2}},Q)$ is a competitive equilibrium. The budget condition is satisfied by $\mathbb{E}^{Q}[X\mathds{1}_{A_{1}}]=\mathbb{E}^{Q}[\xi_{1}]$ and $\mathbb{E}^{Q}[X\mathds{1}_{A_{2}}]=\mathbb{E}^{Q}[X]-\mathbb{E}^{Q}[\xi_{1}]=\mathbb{E}^{Q}[\xi_{2}]$. Market clearance is immediate. It remains to show individual optimality. Denote by $z=\mathbb{E}[u_{1}(X)/X]$. For any $Y$ with $0\leq Y\leq X$ such that $\mathbb{E}^{Q}[Y]\leq\mathbb{E}^{Q}[\xi_{1}]=\mathbb{E}^{Q}[X\mathds{1}_{A_{1}}]$, using that $x\mapsto u_{1}(x)/x$ is increasing, we have
$\mathbb{E}[u_{1}(Y)]=\mathbb{E}\left[Y\frac{u_{1}(Y)}{Y}\right]\leq\mathbb{E}\left[Y\frac{u_{1}(X)}{X}\right]=z\mathbb{E}^{Q}[Y]\leq z\mathbb{E}^{Q}[\xi_{i}]=z\mathbb{E}^{Q}[X\mathds{1}_{A_{1}}]=\mathbb{E}[u_{1}(X_{1})].$
Hence, $\mathbb{E}[u_{1}(Y)]\leq\mathbb{E}[u_{1}(X_{1})]$. For any $Y$ with $0\leq Y\leq X$ such that $\mathbb{E}^{Q}[Y]\leq\mathbb{E}^{Q}[\xi_{2}]=\mathbb{E}^{Q}[X\mathds{1}_{A_{2}}]$, we have
$\mathbb{E}[u_{2}(Y)]\leq\mathbb{E}\left[Y\frac{u_{2}(X)}{X}\right]$ $\leq\mathbb{E}\left[Y\frac{w^{*}u_{1}(X)}{X}\mathds{1}_{A_{1}}\right]+\mathbb{E}\left[Y\frac{u_{2}(X)}{X}\mathds{1}_{A_{2}}\right]$
$=w^{*}z\mathbb{E}^{Q}[Y\mathds{1}_{A_{1}}]+\mathbb{E}\left[Y\frac{u_{2}(X)}{X}\mathds{1}_{A_{2}}\right].$
Moreover, $\mathbb{E}^{Q}[Y]\leq\mathbb{E}^{Q}[X\mathds{1}_{A_{2}}]$ implies $\mathbb{E}^{Q}[Y\mathds{1}_{A_{1}}]\leq\mathbb{E}^{Q}[(X-Y)\mathds{1}_{A_{2}}]$. Hence,
$\mathbb{E}[u_{2}(Y)]$ $\leq w^{*}z\mathbb{E}^{Q}[(X-Y)\mathds{1}_{A_{2}}]+\mathbb{E}\left[Y\frac{u_{2}(X)}{X}\mathds{1}_{A_{2}}\right]$
$\leq\mathbb{E}\left[\frac{X-Y}{X}w^{*}u_{1}(X)\mathds{1}_{A_{2}}\right]+\mathbb{E}\left[Y\frac{u_{2}(X)}{X}\mathds{1}_{A_{2}}\right]$
$\leq\mathbb{E}\left[\frac{X-Y}{X}u_{2}(X)\mathds{1}_{A_{2}}\right]+\mathbb{E}\left[Y\frac{u_{2}(X)}{X}\mathds{1}_{A_{2}}\right]=\mathbb{E}[u_{2}(X\mathds{1}_{A_{2}})].$
Hence, $\mathbb{E}[u_{2}(Y)]\leq\mathbb{E}[u_{2}(X_{2})]$. Therefore, individual optimality holds, and $(X_{1},X_{2},Q)$ is a competitive equilibrium. ∎
B.4 Proofs of results in Section 6
Proof of Theorem 7.
In (i) and (ii) below, since $X$ takes values in $[0,x_{0}]$, we can without loss of generality let $u$ be the identity function. Hence, the preference functional of each agent is $\rho_{h}.$
(i) Write $\mathbf{J}=(\mathds{1}_{A_{1}},\dots,\mathds{1}_{A_{n}})$. First, for each $i\in[n]$, since $h=\overline{h}$ on $[0,1/n]$ and $\mathbb{P}(A_{i})=1/n$, we have
$\rho_{h}(X\mathds{1}_{A_{i}})=\int h(\mathbb{P}(X\mathds{1}_{A_{i}}>x))\,\mathrm{d}x=\int\overline{h}(\mathbb{P}(X\mathds{1}_{A_{i}}>x))\,\mathrm{d}x=\rho_{\overline{h}}(X\mathds{1}_{A_{i}}).$
Moreover,
$\rho_{h}(X\mathds{1}_{A_{i}})=\rho_{\overline{h}}(X\mathds{1}_{A_{i}})=\int\overline{h}\left(\frac{1}{n}\mathbb{P}(X>x)\right)\,\mathrm{d}x.$ (13)
Next, we will show that
$\sum_{i=1}^{n}\rho_{h}(X\mathds{1}_{A_{i}})=\sup_{(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)}\rho_{\overline{h}}(X_{i}),$ (14)
and since $h\leq\overline{h}$, this gives sum-optimality of $(X\mathds{1}_{A_{1}},\dots,X\mathds{1}_{A_{n}})$. To show (14), it suffices to consider jackpot allocations $(X_{1},\dots,X_{n})$ because the preference functional $\rho_{\overline{h}}$ is risk seeking (Chew et al., 1987), and we can apply Theorem 1. Using the representation $(X\mathds{1}_{B_{1}},\dots,X\mathds{1}_{B_{n}})$ of jackpot allocations, we get
$\sup_{(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)}\rho_{\overline{h}}(X_{i})$ $=\sup_{(B_{1},\ldots,B_{n})\in\Pi_{n}}\rho_{\overline{h}}(X\mathds{1}_{B_{i}})$
$=\sup\left\{\sum_{i=1}^{n}\int_{0}^{\infty}\overline{h}(\mathbb{P}(X\mathds{1}_{B_{i}}>x))\mathrm{d}x:(B_{1},\ldots,B_{n})\in\Pi_{n}\right\}$
$\leq\int_{0}^{\infty}\sup\left\{\sum_{i=1}^{n}\overline{h}(p_{i}):\sum_{i=1}^{n}p_{i}=\mathbb{P}(X>x)\right\}\mathrm{d}x$
$=\int_{0}^{\infty}n\overline{h}\left(\frac{1}{n}\mathbb{P}(X>x)\right)\mathrm{d}x,$
where the last equality is due to concavity of $\overline{h}$. Using (13), we get (14), and hence the allocation $(X\mathds{1}_{A_{1}},\dots,X\mathds{1}_{A_{n}})$ is sum-optimal and Pareto optimal.
(ii) Note that since $\overline{h}$ is strictly concave on $(0,1/n)$, $t\mapsto t\overline{h}(x/t)$ is strictly increasing in $t\geq 1$ for each $x\in(0,1/n)$. This implies $n\overline{h}(x/n)>h(x)$ for all $x\in(0,1)$. Hence,
$\frac{1}{n}\int_{0}^{\infty}\overline{h}\left(\mathbb{P}(X>x)\right)\mathrm{d}x<\int_{0}^{\infty}\overline{h}\left(\frac{1}{n}\mathbb{P}(X>x)\right)\mathrm{d}x.$
As a consequence,
$\rho_{h}(X/n)\leq\rho_{\overline{h}}(X/n)=\frac{1}{n}\int_{0}^{\infty}\overline{h}\left(\mathbb{P}(X>x)\right)\mathrm{d}x<\int_{0}^{\infty}\overline{h}\left(\frac{1}{n}\mathbb{P}(X>x)\right)\mathrm{d}x=\rho_{h}(X\mathds{1}_{A_{i}}).$
This shows the strict dominance.
(iii) Note that $h(t/n)/h(t)<1$ for $t\in(0,1)$ because $h$ is concave (thus strictly increasing) on $[0,1/n]$. Together with the continuity of $h$ and condition (10), we get $\sup_{t\in(0,1)}{h(t/n)}/{h(t)}<1$. Hence, we can take $\theta<1$ such that
$\sup_{t\in(0,1)}\frac{h(t/n)}{h(t)}<\theta,$
and take $y_{0}>0$ such that $u(x/n)>\theta u(x)$ for $x\geq y_{0}$, also guaranteed by (10). Note that $X\geq y_{0}$ implies $u(X/n)\geq\theta u(X)$. Hence,
$\rho_{h}(u(X\mathds{1}_{A}))$ $=\int_{0}^{\infty}{h}\left(\mathbb{P}(u(X\mathds{1}_{A})>x)\right)\mathrm{d}x$
$=\int_{0}^{\infty}{h}\left(\frac{\mathbb{P}(u(X)>x)}{n}\right)\mathrm{d}x$
$<\theta\int_{0}^{\infty}{h}(\mathbb{P}(u(X)>x))\mathrm{d}x=\rho_{h}(\theta u(X))\leq\rho_{h}(u(X/n)).$
Therefore, $\overline{\mathbf{X}}$ yields higher utility for each agent than $X\mathbf{J}$ does, and thus strict domination holds.
(iv) Denote by $y=X/n>0$ and $Z^{\varepsilon}$ the first component of $\overline{\mathbf{X}}-\varepsilon\mathbf{1}+n\varepsilon\mathbf{J}$. We can immediately compute
$\rho_{h}(u(Z^{\varepsilon}))=u(y-\varepsilon)+(u(y+(n-1)\varepsilon)-u(y-\varepsilon))h(1/n).$
Taking derivative yields
$\frac{\,\mathrm{d}}{\,\mathrm{d}\varepsilon}\rho_{h}(u(Z^{\varepsilon}))$ $=-u^{\prime}(y-\varepsilon)+((n-1)u^{\prime}(y+(n-1)\varepsilon)+u^{\prime}(y-\varepsilon))h(1/n),$
which converges to $nu^{\prime}(y)(h(1/n)-1/n)$ as $\varepsilon\downarrow 0$. Using $h(1/n)>1/n$, which is implied by the strict concavity of $h$ on $(0,1/n)$, we get that $\,\mathrm{d}\rho_{h}(u(Z^{\varepsilon}))/\,\mathrm{d}\varepsilon>0$ for $\varepsilon>0$ small, and hence $\rho_{h}(u(Z^{\varepsilon}))>u(y)$ for $\varepsilon>0$ small enough. This shows that $\overline{\mathbf{X}}-\varepsilon\mathbf{1}+n\varepsilon\mathbf{J}$ strictly dominates $\overline{\mathbf{X}}$. ∎
Proof of the statement in Example 5.
Let $v=u(y_{0})$. The allocation $\overline{\mathbf{X}}$ yields the maximum utility $v$ to all agents, and hence it is Pareto optimal. To show strict domination, let $(X_{1},\dots,X_{n})$ be a jackpot allocation and $p_{i}=\mathbb{P}(X_{i}>0)$ for $i\in[n]$. With this allocation, agent $i$ has utility
$\int_{0}^{v}\overline{h}(\mathbb{P}(u(X_{i})>x))\,\mathrm{d}x\leq vh(p_{i}).$
Since at least one $p_{i}$ is less than or equal to $1/n$, the condition $h(1/n)<1$ guarantees that at least one agent has a utility less than $v$, and thus the jackpot allocation is strictly dominated by $\overline{\mathbf{X}}$. ∎
Proof of Proposition 5.
Without loss of generality, assume that $u$ is the identity on $[0,x_{0}]$. The budget conditions and market clearance hold by construction of $x\mathbf{J}$. We only need to show individual optimality. Write $(\mathds{1}_{A_{1}},\dots,\mathds{1}_{A_{n}})=\mathbf{J}$. Note that for any random variable $Y$ taking values in $[0,1]$ we have $Y\leq_{\rm cx}\mathds{1}_{A}$ where $\mathbb{P}(A)=\mathbb{E}[Y]$. Therefore, we have $\rho_{\overline{h}}(Y)\leq\rho_{\overline{h}}(\mathds{1}_{A})=\overline{h}(\mathbb{E}[Y])$ because $\rho_{\overline{h}}$ represents a risk-seeking preference. It follows that for any random variable $X_{i}$ in $[0,x]$ satisfying the constraint $\mathbb{E}[X_{i}]=\mathbb{E}[\xi_{i}]$, we have
$\rho_{h}(X_{i})\leq\rho_{\overline{h}}(X_{i})\leq\rho_{\overline{h}}(x\mathds{1}_{A_{i}})=x\overline{h}(\mathbb{P}(A_{i}))=xh(\mathbb{P}(A_{i}))=\rho_{h}(x\mathds{1}_{A_{i}}).$
This extends to the case $\mathbb{E}[X_{i}]<\mathbb{E}[\xi_{i}]$ by monotonicity. Therefore, individual optimality holds, showing that $(x\mathbf{J},\mathbb{P})$ is a competitive equilibrium. ∎
Appendix C Technical background on counter-monotonicity
We provide some technical background on counter-monotonicity. First, Dall’Aglio (1972) obtained some necessary conditions for counter-monotonicity in dimensions more than two. As standard in probability theory, a random variable is non-degenerate if it is not almost surely a constant.
Proposition 6 (Dall’Aglio (1972)).
If at least three of $X_{1},\dots,X_{n}$ are non-degenerate, counter-monotonicity of $(X_{1},\dots,X_{n})$ is equivalent to that one of the following two cases holds true:
$\mathbb{P}(X_{i}>\operatorname*{ess\text{-}inf}X_{i},~{}X_{j}>\operatorname*{ess\text{-}inf}X_{j})=0\mbox{ for all $i\neq j$};$ (15)
$\mathbb{P}(X_{i}<\operatorname*{ess\text{-}sup}X_{i},~{}X_{j}<\operatorname*{ess\text{-}sup}X_{j})=0\mbox{ for all $i\neq j$}.$ (16)
A necessary condition for (15) is $\sum_{i=1}^{n}\mathbb{P}(X_{i}>\operatorname*{ess\text{-}inf}X_{i})\leq 1,$ and a necessary condition for (16) is $\sum_{i=1}^{n}\mathbb{P}(X_{i}<\operatorname*{ess\text{-}sup}X_{i})\leq 1.$
The alternative formulation (3) of jackpot allocations in Section 3 directly follows from Proposition 6.
Recall that $\Pi_{n}$ is the set of all $n$-compositions of $\Omega$. The next proposition, which is a reformulation of Lauzier et al. (2023a, Theorem 1), simplifies the stochastic representation of counter-monotonicity. Note that the equalities below, as well as other equalities and inequalities between random variables, are understood to hold almost surely. In the next result, $\mathcal{X}=L^{1}$.
Proposition 7.
For $X\in\mathcal{X}$, suppose that at least three of $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ are non-degenerate. Then, $(X_{1},\dots,X_{n})$ is counter-monotonic if and only if there exist constants $m_{1},\dots,m_{n}$ and $(A_{1},\dots,A_{n})\in\Pi_{n}$ such that
either $X_{i}=(X-m)\mathds{1}_{A_{i}}+m_{i}~{}~{}~{}\mbox{for all $i\in[n]$ with}~{}m=\sum_{i=1}^{n}m_{i}\leq\operatorname*{ess\text{-}inf}X;$ (17)
or $X_{i}=(X-m)\mathds{1}_{A_{i}}+m_{i}~{}~{}~{}\mbox{for all $i\in[n]$ with}~{}m=\sum_{i=1}^{n}m_{i}\geq\operatorname*{ess\text{-}sup}X.$ (18)
Proof of Proposition 7.
The “if” part follows from the fact that $\sum_{i=1}^{n}X_{i}=X$ and Proposition 6. We will show the “only if” part. Assume that $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is counter-monotonic. By Lauzier et al. (2023a, Theorem 1), there exists $(A_{1},\dots,A_{n})\in\Pi_{n}$ such that
$X_{i}=(X-m)\mathds{1}_{A_{i}}+m_{i}~{}~{}~{}\mbox{for all $i\in[n]$},$
where either $m_{i}=\operatorname*{ess\text{-}inf}X_{i}$ for all $i\in[n]$ or $m_{i}=\operatorname*{ess\text{-}sup}X_{i}$ for all $i\in[n]$, and $m=\sum_{i=1}^{n}m_{i}$. If $m_{i}=\operatorname*{ess\text{-}inf}X_{i}$ for all $i\in[n]$, we have $m=\sum_{i=1}^{n}\operatorname*{ess\text{-}inf}(X_{i})\leq\operatorname*{ess\text{-}inf}(\sum_{i=1}^{n}X_{i})=\operatorname*{ess\text{-}inf}X$. If $m_{i}=\operatorname*{ess\text{-}sup}X_{i}$ for all $i\in[n]$, we have $m=\sum_{i=1}^{n}\operatorname*{ess\text{-}sup}(X_{i})\geq\operatorname*{ess\text{-}sup}(\sum_{i=1}^{n}X_{i})=\operatorname*{ess\text{-}sup}X$. ∎
The allocation $(X,0,\dots,0)$ is counter-monotonic by taking $A=\Omega$ and $m=m_{1}=\operatorname*{ess\text{-}inf}X$, and it is trivial to verify that it is also comonotonic. Notice now that the allocations defined in equation (17) and (18) reflect the conditions in Proposition 6. In (17), for almost every $\omega\in\Omega$, at most one agent receives more than their essential infimum. Conversely, in (18), at most one agent receives less than their essential supremum.
Appendix D Some extensions of Theorem 7
We briefly discuss a few dimensions in which the statements in Theorem 7 can readily be generalized. We did not pursue these generalizations because they do not seem to offer stronger empirical relevance than Assumption H-RDU.
  1. (a)
    Assumption H-RDU allows for $h$ to be concave. In this case, $\overline{h}=h$, and the agents are risk seeking for payoffs valued in $[0,x_{0}]$.
  2. (b)
    By inspecting the proof of Theorem 7, it suffices to require $h=\overline{h}$ on $[0,1/n]$, and whether $h$ is concave-convex beyond $1/n$ is irrelevant.
  3. (c)
    The result remains true if $u$ is convex on $[0,x_{0}]$ instead of being linear, following the same proof, by noting that an agent with a convex utility function and the probability weighting function $\overline{h}$ is risk seeking, which is the main step to apply Theorem 1.
  4. (d)
    A careful inspection of the proofs of main results reveals that, for most of our results on risk-seeking EU agents, it suffices to assume that $x\mapsto u(x)/x$ is increasing instead of the convexity of $u$ (this condition is weaker than convexity with $u(0)=0$). Moreover, for the RDU agents in Assumption H-RDU, we can use this condition on $[0,x_{0}]$ instead of linearity, and the results in Theorem 7 hold true.
We formally prove the assertion in (d) below. Let $u$ be an increasing function with $u(0)=0$ and $x\mapsto u(x)/x$ is increasing. We first show an analogue of Theorem 1. Let $(X_{1},\dots,X_{n})$ and $(Y_{1},\dots,Y_{n})$ be as in the proof of Theorem 1. Note that
$u(Y_{i})=u\left(X\mathds{1}_{\{Z_{i-1}\leq U<Z_{i}\}}\right)=u(X)\mathds{1}_{\{Z_{i-1}\leq U<Z_{i}\}}\geq\frac{Xu(X_{i})}{X_{i}}\mathds{1}_{\{Z_{i-1}\leq U<Z_{i}\}}\mathds{1}_{\{X_{i}>0\}}.$
Hence,
$\mathbb{E}\left[u(Y_{i})\mid X_{1},\dots,X_{n}\right]$ $\geq\frac{Xu(X_{i})}{X_{i}}\mathds{1}_{\{X_{i}>0\}}\mathbb{E}\left[\mathds{1}_{\{Z_{i-1}\leq U<Z_{i}\}}\mid X_{1},\dots,X_{n}\right]$
$=\frac{Xu(X_{i})}{X_{i}}\mathds{1}_{\{X_{i}>0\}}\frac{X_{i}}{X}$
$=u(X_{i})\mathds{1}_{\{X_{i}>0\}}=u(X_{i}).$
This shows $u(Y_{i})\geq_{\rm icx}u(X_{i})$, where $\geq_{\rm icx}$ is increasing convex order (meaning $\mathbb{E}[\phi(Y_{i})]\geq\mathbb{E}[\phi(X_{i})]$ for all increasing convex $\phi$). This implies $\rho_{\overline{h}}(u(Y_{i}))\geq\rho_{\overline{h}}(u(X_{i}))$ because $\rho_{\overline{h}}$ is increasing in convex order (e.g., Wang et al. 2020, Theorem 3), and increasing convex order can be decomposed into convex order and first-order stochastic dominance (e.g., Shaked and Shanthikumar 2007, Theorem 4.A.6). Therefore, the jackpot allocation $(Y_{1},\dots,Y_{n})$ dominates $(X_{1},\dots,X_{n})$, and for sum-optimality it suffices to consider jackpot allocations. The rest of the proof follows the same arguments in the proof of parts (i) and (ii) of Theorem 7 with $X$ replaced by $u(X)$. Parts (iii) and (iv) do not rely on the properties of $u$ on $[0,x_{0}]$.
We finally note that if $u$ is convex on $[0,a]$ and concave on $[a,\infty)$, then $x\mapsto u(x)/x$ is increasing on some interval $[0,b]$ with $b\geq a$ (often $b>a$).
Appendix E On the existence of competitive equilibria for EU agents
For the interested reader, we discuss the existence of competitive equilibria in the setting of Section 5 (under Assumption EURS) for a given initial endowment $(\xi_{1},\dots,\xi_{n})\in\mathbb{A}_{n}(X)$. We outline a general approach under some assumptions, and prove the existence of competitive equilibria in the special case of proportional endowments.
First, we make an assumption of no-ties in the weighted utility functions.
Assumption NT.
For $i\neq j$, $\{x\in\mathbb{R}_{+}:\lambda_{i}u_{i}(x)=\lambda_{j}u_{j}(x))\}$ is finite for any $\lambda_{i},\lambda_{j}>0$, and $X$ is continuously distributed.
A simple example of utility functions satisfying Assumption NT is that agent $i$ is more risk seeking than agent $i+1$ for $i\in[n]$; that is, $u_{i}=T_{i}\circ u_{i+1}$ for some increasing and strictly convex function $T_{i}$ for $i\in[n-1]$. In this case, for any $\lambda_{i},\lambda_{j}>0$, $i\neq j$, we have that $\lambda_{i}u_{i}$ and $\lambda_{j}u_{j}$ cross at most once under Assumption EURS. For instance, this holds for $u_{i}(x)=x^{\alpha_{i}}$, $i\in[n]$, with distinct values of $\alpha_{i}$. For an illustration, see the left panel of Figure 2.
For any given ${\bm{\lambda}}=(\lambda_{1},\dots,\lambda_{n})\in\Delta_{n}$, define the sets $A^{\bm{\lambda}}_{i}=\{\lambda_{i}u_{i}(X)=V_{\bm{\lambda}}(X)\}$ for $i\in[n]$ and let $\mathbf{J}^{{\bm{\lambda}}}=(\mathds{1}_{A^{{\bm{\lambda}}}_{1}},\dots,\mathds{1}_{A^{{\bm{\lambda}}}_{n}})$. For $i\neq j$, Assumption NT implies that $\mathbb{P}(\lambda_{i}u_{i}(X)=\lambda_{j}u_{j}(X))=0$ because $X$ is continuously distributed. Hence, $\mathbb{P}(A_{i}^{{\bm{\lambda}}}\cap A_{j}^{{\bm{\lambda}}})=0$ and $\sum_{i=1}^{n}\mathds{1}_{A^{\bm{\lambda}}_{i}}=1$ (almost surely). Thus, $X\mathbf{J}^{{\bm{\lambda}}}$ is a jackpot allocation.
Next, we introduce three useful objects. Define a probability measure $Q^{\bm{\lambda}}$ by
$\frac{\,\mathrm{d}Q^{\bm{\lambda}}}{\,\mathrm{d}\mathbb{P}}=\frac{V_{\bm{\lambda}}(X)}{X}\frac{1}{\mathbb{E}[V_{\bm{\lambda}}(X)/X]}\mbox{ ~{}~{} with the convention $0/0=0$},$
a function $f:\Delta_{n}\to\Delta_{n}$ by
$f({\bm{\lambda}})=\left(\frac{\mathbb{E}^{Q^{\bm{\lambda}}}[X\mathds{1}_{A^{\bm{\lambda}}_{1}}]}{\mathbb{E}^{Q^{\bm{\lambda}}}[X]},\dots,\frac{\mathbb{E}^{Q^{\bm{\lambda}}}[X\mathds{1}_{A^{\bm{\lambda}}_{n}}]}{\mathbb{E}^{Q^{\bm{\lambda}}}[X]}\right)=\frac{\mathbb{E}^{Q^{\bm{\lambda}}}[X\mathbf{J}^{{\bm{\lambda}}}]}{\mathbb{E}^{Q^{\bm{\lambda}}}[X]},~{}~{}~{}{\bm{\lambda}}\in\Delta_{n},$
and a function $g:\Delta_{n}\to\Delta_{n}$ by
$g({\bm{\lambda}})=\left(\frac{\mathbb{E}^{Q^{\bm{\lambda}}}[\xi_{1}]}{\mathbb{E}^{Q^{\bm{\lambda}}}[X]},\dots,\frac{\mathbb{E}^{Q^{\bm{\lambda}}}[\xi_{n}]}{\mathbb{E}^{Q^{\bm{\lambda}}}[X]}\right),~{}~{}~{}{\bm{\lambda}}\in\Delta_{n}.$
Note that $\mathbb{E}^{Q^{\bm{\lambda}}}[X]>0$ holds under Assumption EURS. Our goal is to find ${\bm{\lambda}}\in\Delta_{n}$ such that $f({\bm{\lambda}})=g({\bm{\lambda}})$. The next proposition justifies that finding such ${\bm{\lambda}}$ is sufficient for finding a competitive equilibrium.
Proposition 8.
Under Assumptions EURS and NT, if ${\bm{\lambda}}\in\Delta_{n}$ and $f({\bm{\lambda}})=g({\bm{\lambda}})$, then $(X\mathbf{J}^{{\bm{\lambda}}},Q^{{\bm{\lambda}}})$ is a competitive equilibrium for the vector of initial endowments $(\xi_{1},\dots,\xi_{n})\in\mathbb{A}_{n}(X)$.
Proof of Proposition 8.
Write $(X_{1},\dots,X_{n},Q)=(X\mathbf{J}^{{\bm{\lambda}}},Q^{{\bm{\lambda}}})$ and $(\lambda_{1},\dots,\lambda_{n})={\bm{\lambda}}$. The equality $f({\bm{\lambda}})=g({\bm{\lambda}})$ implies $\mathbb{E}^{Q}[X_{i}]=\mathbb{E}^{Q}[\xi_{i}]$ for $i\in[n]$. Let $z=\mathbb{E}[V_{\bm{\lambda}}(X)/X]$. Fix $i\in[n]$. For any $Y$ with $0\leq Y\leq X$ such that $\mathbb{E}^{Q}[Y]\leq\mathbb{E}^{Q}[\xi_{i}]$, we have
$\mathbb{E}[\lambda_{i}u_{i}(Y)]=\mathbb{E}\left[Y\frac{\lambda_{i}u_{i}(Y)}{Y}\right]\leq\mathbb{E}\left[Y\frac{\lambda_{i}u_{i}(X)}{X}\right]\leq\mathbb{E}\left[Y\frac{V_{\bm{\lambda}}(X)}{X}\right]=z\mathbb{E}^{Q}[Y]\leq z\mathbb{E}^{Q}[\xi_{i}]=z\mathbb{E}^{Q}[X_{i}].$
Moreover, since $A^{\bm{\lambda}}_{i}=\{\lambda_{i}u_{i}(X)=V_{\bm{\lambda}}(X)\}$, we have $V_{\bm{\lambda}}(X)\mathds{1}_{A^{\bm{\lambda}}_{i}}=\lambda_{i}u_{i}(X)\mathds{1}_{A^{\bm{\lambda}}_{i}}=\lambda_{i}u_{i}(X_{i})$, and this implies
$z\mathbb{E}^{Q}[X_{i}]=\mathbb{E}\left[X\mathds{1}_{A^{\bm{\lambda}}_{i}}\frac{V_{\bm{\lambda}}(X)}{X}\right]=\mathbb{E}[\lambda_{i}u_{i}(X_{i})].$
Hence, $\mathbb{E}[u_{i}(Y_{i})]\leq\mathbb{E}[u_{i}(X_{i})]$ and thus $X_{i}$ satisfies individual optimality for agent $i$. The market clearance condition $\sum_{i=1}^{n}X_{i}=X$ holds true because $\sum_{i=1}^{n}\mathds{1}_{A^{\bm{\lambda}}_{i}}=1$. Therefore, $(X_{1},\dots,X_{n},Q)$ is a competitive equilibrium. ∎
The remaining task is to find ${\bm{\lambda}}$ with $f({\bm{\lambda}})=g({\bm{\lambda}})$. We do not know a general solution to this problem, but in the simplified scenario of proportional endowments, the problem can be solved.
Assumption PE.
The initial endowment vector $(\xi_{1},\dots,\xi_{n})$ is equal to $(\theta_{1}X,\dots,\theta_{n}X)$ for some $(\theta_{1},\dots,\theta_{n})\in\Delta_{n}$.
Under Assumption PE, we have $g({\bm{\lambda}})=(\theta_{1},\dots,\theta_{n})$ for any ${\bm{\lambda}}\in\Delta_{n}$. In this situation, we can show that $f({\bm{\lambda}})=g({\bm{\lambda}})$ holds, through a technique established by Jamison and Ruckle (1976).
Proposition 9.
If Assumptions EURS, NT and PE hold, then there exists a competitive equilibrium of the form $(X\mathbf{J}^{{\bm{\lambda}}},Q^{{\bm{\lambda}}})$ for some ${\bm{\lambda}}\in\Delta_{n}$.
Proof of Proposition 9.
A face of $\Delta_{n}$ is the set $\Delta_{n}^{D}=\{(x_{1},\dots,x_{n})\in\Delta_{n}:x_{j}=0~{}\mbox{for}~{}j\in D\}$ for some $D\subseteq[n]$. Lemma 6 below guarantees that $f$ is a continuous function that carries each face of $\Delta_{n}$ into itself. This condition allows us to apply Jamison and Ruckle (1976, Lemma 2.1), which implies that $f$ is surjective. Hence, there exists ${\bm{\lambda}}\in\Delta_{n}$ such that $f({\bm{\lambda}})=(\theta_{1},\dots,\theta_{n})=g({\bm{\lambda}})$. By Proposition 8, $(X\mathbf{J}^{{\bm{\lambda}}},Q^{\bm{\lambda}})$ is a competitive equilibrium. ∎
Lemma 6.
If Assumptions EURS and NT hold, then $f$ is a continuous function that carries each face of $\Delta_{n}$ into itself.
Proof of Lemma 6.
For $i\in[n]$, define $f_{i}:\Delta_{n}\to\mathbb{R}_{+}$ by $f_{i}({\bm{\lambda}})=\mathbb{E}^{Q^{\bm{\lambda}}}[X\mathds{1}_{A^{\bm{\lambda}}_{i}}]$ for ${\bm{\lambda}}\in\Delta_{n}$. We have
$f_{i}({\bm{\lambda}})=\mathbb{E}^{Q^{\bm{\lambda}}}[X\mathds{1}_{A^{\bm{\lambda}}_{i}}]=\mathbb{E}[V_{\bm{\lambda}}(X)\mathds{1}_{A^{\bm{\lambda}}_{i}}]=\mathbb{E}[\lambda_{i}u_{i}(X)\mathds{1}_{A^{\bm{\lambda}}_{i}}].$
Let ${\bm{\lambda}}=(\lambda_{1},\dots,\lambda_{n})$ and ${\bm{\zeta}}=(\zeta_{1},\dots,\zeta_{n})\in\Delta_{n}$ be such that $\|{\bm{\lambda}}-{\bm{\zeta}}\|:=\sum_{i=1}^{n}|\lambda_{i}-\zeta_{i}|<\varepsilon$. As $X$ is continuously distributed and $u_{1},\dots,u_{n}$ are continuous, Assumption NT implies
$p:=\mathbb{P}(A_{i}^{\bm{\lambda}}\cup A_{i}^{\bm{\zeta}})-\mathbb{P}(A_{i}^{\bm{\lambda}}\cap A_{i}^{\bm{\zeta}})=\mathbb{P}\left(X\in\bigcup_{x^{*}:\lambda_{i}u_{i}(x^{*})=\lambda_{j}u_{j}(x^{*})}\{x^{*}-c_{1}\varepsilon<x<x^{*}+c_{1}\varepsilon\}\right)<c_{2}\varepsilon$
for some $c_{1},c_{2}>0$, because switching from $A_{i}^{\bm{\lambda}}$ to $A_{i}^{\bm{\zeta}}$ or back can only happen at points in some neighborhoods of $\{x:\lambda_{i}u_{i}(x)=\lambda_{j}u_{j}(x)\}$. This implies that ${\bm{\lambda}}\mapsto\mathbb{E}[u_{i}(X)\mathds{1}_{A_{i}^{{\bm{\lambda}}}}]$ is continuous, further guaranteeing that $f_{i}$ is continuous. Therefore, $\hat{f}:=\sum_{i=1}^{n}f_{i}$ is also a continuous function. Moreover, under Assumption EURS, we have $\hat{f}>0$. Since $f=(f_{1}/\hat{f},\dots,f_{n}/\hat{f})$, we know that $f$ is continuous. Moreover, if $\lambda_{i}=0$, then $\mathbb{P}(A^{\bm{\lambda}}_{i})=0$ and thus $\mathbb{E}^{Q}[X\mathds{1}_{A^{\bm{\lambda}}_{i}}]/\mathbb{E}^{Q}[X]=0$. Hence, for any face $\Delta_{n}^{D}$ and ${\bm{\lambda}}\in\Delta_{n}^{D}$, we have $f({\bm{\lambda}})\in\Delta_{n}^{D}$. ∎